### v3

# Does FLUX Already Know How to Perform Physically Plausible Image Composition?

Abstract
Image composition aims to seamlessly insert a user-specified object into a new scene, but existing models struggle with complex lighting (e.g., accurate shadows, water reflections) and diverse, high-resolution inputs. Modern text-to-image diffusion models (e.g., SD3.5, FLUX) already encode essential physical and resolution priors, yet lack a framework to unleash them without resorting to latent inversion, which often locks object poses into contextually inappropriate orientations, or brittle attention surgery. We propose SHINE, a training-free framework for Seamless, High-fidelity Insertion with Neutralized Errors. SHINE introduces manifold-steered anchor loss, leveraging pretrained customization adapters (e.g., IP-Adapter) to guide latents for faithful subject representation while preserving background integrity. Degradation-suppression guidance and adaptive background blending are proposed to further eliminate low-quality outputs and visible seams. To address the lack of rigorous benchmarks, we introduce ComplexCompo, featuring diverse resolutions and challenging conditions such as low lighting, strong illumination, intricate shadows, and reflective surfaces. Experiments on ComplexCompo and DreamEditBench show state-of-the-art performance on standard metrics (e.g., DINOv2) and human-aligned scores (e.g., DreamSim, ImageReward, VisionReward). Code and benchmark will be publicly available upon publication.
Figure 1: Showcase of our training-free image composition method, SHINE. This gallery highlights SHINE’s ability to seamlessly integrate subjects into complex scenes, including low-light conditions, intricate shadows, and water reflections.
1 Introduction
Image composition, which places a user–specified object into a new scene, is a demanding image editing task. Despite the breathtaking progress of multimodal foundation models (e.g., GPT-5 (OpenAI, 2025), Gemini-2.5 (Gemini2.5, 2025), SeedEdit/Doubao (Shi et al., 2024b), and Grok-4 (gro, 2025)), these generic models still struggle with image composition. Typical failures include imprecise object placement, inconsistent lighting, and the subject’s identity drift (see Fig. 2). These limitations indicate that, as of now, massive multimodal pre-training alone has not yet endowed them with sufficient compositional ability for this task. A natural response has been to train specialized models. Yet building large-scale, high-quality, multi-resolution triplet datasets (object, scene, composite) is prohibitively costly. As a result, most composition models are fine-tuned from base models (e.g., FLUX.1-dev (Black Forest Labs, 2024a), FLUX.1-Fill (Black Forest Labs, 2024c), SDXL (Podell et al., 2024)) on synthetic data generated via inpainting or augmentations (Chen et al., 2024c; Yang et al., 2023; Song et al., 2023; Wang et al., 2025; He et al., 2024).
These models, however, face two main limitations (see Fig. 6): (i) Lighting realism. They struggle to achieve natural composition under complex lighting conditions, such as accurate shadow generation or water reflections for the inserted subject. (ii) Resolution rigidity. They are tied to a fixed resolution, necessitating downsampling or cropping when applied to varied, high-resolution background images, which degrades generation quality. Notably, such issues are absent in the base models, implying that the underlying physical priors are present but are not effectively exploited by fine-tuned variants. The degradation largely stems from low-quality synthetic datasets, which inherit flaws from inpainting models that often mis-handle shadows and reflections, producing implausible edits, hallucinated content, or incomplete object removal (Yu et al., 2025b; Winter et al., 2024).
There have been prior training-free attempts to exploit the priors of text-to-image (T2I) models for advancing image composition, but they fall short for two main reasons. (i) Inversion bottlenecks. Most methods (Lu et al., 2023d; Pham et al., 2024; Yan et al., 2025; Li et al., 2024b) depend on accurate image inversion (Song et al., 2020; Lu et al., 2022; Mokady et al., 2023). In practice, inversion constrains the inserted object to the pose of its reference image, often resulting in contextually inappropriate orientations. Moreover, inversion is less effective for classifier-free guidance (CFG) distilled models (e.g., FLUX), where elevated inversion errors degrade identity preservation. (ii) Fragile attention surgery. Many training-free approaches rely on attention manipulation (Lu et al., 2023d; Yan et al., 2025; Li et al., 2024b). While compatible with the joint self-attention in Multimodal Diffusion Transformers (MMDiT) (Peebles & Xie, 2023), these methods inherit the instability and hyperparameter sensitivity (Lu et al., 2023d), limiting their robustness.
To bridge these gaps we present SHINE, a training-free framework for Seamless, High-fidelity Insertion with Neutralized Errors (see Fig. 1). SHINE comprises three innovations: (i) Manifold-Steered Anchor (MSA) loss, which leverages pretrained open-domain customization adapters (e.g., IP-Adapter (Ye et al., 2023b)) to steer noisy latents toward faithfully representing the reference subject while preserving the structural integrity of the background. (ii) Degradation-Suppression Guidance (DSG) that steers sampling away from low-quality distributions. (iii) Adaptive Background Blending (ABB) that eliminates visible seams along mask boundaries.
Existing benchmarks primarily comprise background images with a fixed resolution of $512\times 512$ pixels. To evaluate performance across diverse, high-resolution, and demanding scenarios, we introduce ComplexCompo, a benchmark that includes varied resolutions, both landscape and portrait orientations, and complex conditions such as low lighting, intense illumination, intricate shadows, and water reflections. Extensive experiments on ComplexCompo and DreamEditBench (Li et al., 2023b) demonstrate that SHINE achieves state-of-the-art (SOTA) performance, surpassing baselines on standard metrics (e.g., DINOv2 (Oquab et al., 2023)) and human-aligned metrics (e.g., DreamSim (Fu et al., 2023), ImageReward (Xu et al., 2023), VisionReward (Xu et al., 2024)).
Figure 2: Image composition from advanced multimodal models under three challenging conditions: backlighting, shadows, and water surfaces. Refer to Appendix H for prompt details.
2 Related Work
This section reviews prior work on image composition. A more comprehensive discussion, covering image composition, general image editing, and subject-driven generation, is offered in Appendix A. Classical image composition splits into sub-tasks (Niu et al., 2021) such as object placement (Azadi et al., 2020; Zhang et al., 2020a), blending (Wu et al., 2019; Zhang et al., 2020b), harmonization (Cao et al., 2023; Lu et al., 2023b), and shadow generation (Hong et al., 2022; Sheng et al., 2021), typically handled by separate models. Diffusion models have shifted the field toward unified frameworks, either training-based or training-free. Training-based approaches fine-tune diffusion models with curated datasets, adding grounding layers, controllability signals, or identity-preserving supervision from image or video sets (Wang et al., 2025; Chen et al., 2024c; Yang et al., 2023; Song et al., 2023; Lu et al., 2023c). However, they often bias model priors and struggle with complex lighting due to the lack of large-scale real-world triplets. Training-free approaches avoid retraining by manipulating inversion and attention during inference, enabling flexible test-time adaptation (Yan et al., 2025; Li et al., 2024b; 2023b; Lu et al., 2023d; Pham et al., 2024). Yet these methods remain fragile: strong injections preserve identity but fix unnatural poses, while weaker ones improve realism at the cost of fidelity, reflecting a core trade-off between identity preservation and natural composition.
3 Method
Image composition seeks to integrate a subject into a designated area of a background image while preserving the integrity of the surrounding scene. This process typically requires three inputs: (1) one or more reference images of the subject $\{\bm{x}^{\text{subj}}_{1},\bm{x}^{\text{subj}}_{2},\ldots,\bm{x}^{\text{subj}}_{n}\}$, (2) a background image $\bm{x}^{\text{bg}}$, and (3) a user-provided mask $\bm{M}^{\text{user}}$ specifying the insertion region within the background.
Our framework is built on three core components: Manifold-Steered Anchor (MSA) loss, Degradation-Suppression Guidance (DSG), and Adaptive Background Blending (ABB). Importantly, the design is model-agnostic and requires only standard features of modern generative models: MSA loss assumes that the base model supports either personalization finetuning or provides access to a pretrained personalization adapter, DSG uses self-attention maps, and ABB relies on text-image cross-attention. These mild assumptions enable seamless integration into existing pipelines without architectural changes. We present main results with FLUX, while additional experiments on SDXL (Podell et al., 2024), SD3.5 (Esser et al., 2024), and PixArt (Chen et al., 2023) are provided in Appendix E. The complete algorithm is shown in Algorithm 1.
Figure 3: Overview of the proposed framework. (a) The noisy latent is created by inpainting the background with a VLM-derived object description, then adding Gaussian noise. (b) Manifold-Steered Anchor (MSA) loss guides noisy latents toward faithfully capturing the reference subject (red arrow), while preserving the structural integrity of the background. Concretely, it enforces that the prediction of the optimized latent $\bm{z}_{t}^{*}$ on the adapter-augmented model’s manifold remains close to the prediction of the original latent $\bm{z}_{t}$ on the base model’s manifold. (c) Degradation-Suppression Guidance (DSG) constructs a negative velocity pointing toward low-quality regions by blurring $\bm{Q}_{\text{img}}$ and, in a CFG-like manner, steers the trajectory away from this low-quality distribution.
3.1 Non-Inversion Latent Preparation
In training-free diffusion-based image composition (Lu et al., 2023d; Pham et al., 2024; Yan et al., 2025; Li et al., 2024b), it is common to start from a noisy latent. Existing training-free frameworks typically rely on image inversion, where the initial noisy latent is constructed by copying the inverted latent of the subject image into a designated region of the background image’s inverted latent.
However, this copy-paste strategy constrains the inserted object to the exact pose of its reference image, often leading to contextually inappropriate orientations in the composed result. Moreover, inversion is suboptimal for CFG-distilled models (e.g., FLUX), as it introduces higher inversion errors that compromise subject identity preservation.
To address these limitations, we abandon inversion and instead perform a one-step forward diffusion to obtain the noisy latent. As illustrated in Fig. 3(a), we use a vision-language model (VLM) (Xue et al., 2024; Chen et al., 2024d; Liu et al., 2024) to caption the subject image and leverage this caption, along with an image inpainting model (Li et al., 2024c; Ju et al., 2024; Zhuang et al., 2024; Black Forest Labs, 2024b), to generate the inpainted background image, denoted as $\bm{x}^{\text{init}}$. The noisy latent is encoded in the VAE space as $\bm{z}^{\text{init}}$ and perturbed to timestep $t\leq T$ via one-step forward diffusion, following the flow matching formulation: $\bm{z}_{t}=(1-\sigma_{t})\bm{z}^{\text{init}}+\sigma_{t}\bm{\epsilon}$, where $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{1})$.
3.2 Manifold-Steered Anchor Loss
The Manifold-Steered Anchor (MSA) loss is designed to optimize the noisy latent $\bm{z}_{t}$ (from Sec. 3.1) during the denoising process, steering it toward a reference subject while preserving the structural integrity of the original image. The key intuition is to leverage the prior knowledge embedded in pretrained open-domain customization adapters (or alternatively, personalized LoRAs) such as IP-Adapter (Ye et al., 2023b), PuLID (Guo et al., 2024), and InstantCharacter (Tao et al., 2025), to intervene directly in the diffusion trajectory. Specifically, the MSA loss is defined as:
$\mathop{\min}_{\bm{z}_{t}}\mathcal{L}_{\text{MSA}}(\bm{z}_{t})=\Big\|\bm{v}_{\bm{\theta}+\bm{\Delta\theta}}(\bm{z}_{t},t,\bm{c},\bm{z}^{\text{subj}})-{\tt sg}[\tilde{\bm{v}}_{t}]\Big\|_{2}^{2},$ (1)
where $\tilde{\bm{v}}_{t}\triangleq\bm{v}_{\bm{\theta}}(\tilde{\bm{z}}_{t},t,\bm{c})$ serves as a fixed anchor, preserving the structure of the background image at a given noise level $t$, with $\tilde{\bm{z}}_{t}$ held constant as the original noisy latent. $\bm{v}_{\bm{\theta}}(\cdot)$ denotes the velocity predicted by the frozen T2I model $\bm{\theta}$, while $\bm{v}_{\bm{\theta}+\bm{\Delta\theta}}(\cdot)$ represents the velocity predicted by the a T2I model augmented with an adapter $\Delta\bm{\theta}$. $\bm{z}^{\text{subj}}$ is the latent of the subject image. The text prompt $\bm{c}$ is from the VLM’s description of $\bm{x}^{\text{init}}$, and $\tt sg$$\cdot$] indicates the stop-gradient operation.
The MSA loss is motivated by the observation that optimizing a latent representation against a frozen generative model implicitly projects the latent onto the model’s learned data manifold (Meng et al., 2021; Kim et al., 2022; Graikos et al., 2022; Feng et al., 2023). The generator serves as an implicit prior, guiding gradient descent toward the manifold’s basin of attraction.
For instance, when a generative model $\bm{G}(\bm{w})$ is trained solely on cat images, its outputs are confined to the cat-image manifold. Thus, approximating a dog image $\bm{x}_{\text{dog}}$ by solving $\mathop{\min}_{\bm{w}}\|\bm{G}(\bm{w})-\bm{x}_{\text{dog}}\|_{2}^{2}$ yields $\bm{G}(\bm{w}^{*})$ that remains a cat image, but with structural features (e.g., pose or outline) aligned to $\bm{x}_{\text{dog}}$. The result is the projection of the dog image onto the cat manifold, not a genuine dog image.
Analogously, MSA loss is designed to achieve two goals simultaneously. (1) It seeks an optimized noisy latent $\bm{z}_{t}^{*}$ that remains within the manifold of the adapter-augmented model when conditioned on the subject $\bm{z}^{\text{subj}}$. (2) It encourages the adapter’s prediction on this latent $\bm{z}_{t}^{*}$ to align with the base model’s prediction on the original latent $\bm{z}_{t}$, i.e., $\bm{v}_{\bm{\theta}+\bm{\Delta\theta}}(\bm{z}_{t}^{*},t,\bm{c},\bm{z}^{\text{subj}})\approx\bm{v}_{\bm{\theta}}(\bm{z}_{t},t,\bm{c})$ (see Fig. 3(b)). Since the velocity prediction of a T2I model on a noisy latent $\bm{z}_{t}$ can also be interpreted as a coarse estimate of the clean image that encodes essential structural information (Zheng et al., 2023), this alignment preserves the spatial layout and background details inherited from the original image.
The gradient of $\mathcal{L}_{\text{MSA}}$ with respect to $\bm{z}_{t}$ is:
$\nabla_{\bm{z}_{t}}\mathcal{L}_{\text{MSA}}(\bm{z}_{t})=2\Big(\bm{v}_{\bm{\theta}+\bm{\Delta\theta}}(\bm{z}_{t},t,\bm{c},\bm{z}^{\text{subj}})-\text{sg}[\tilde{\bm{v}}_{t}]\Big)\frac{\partial\bm{v}_{\bm{\theta}+\bm{\Delta\theta}}(\bm{z}_{t},t,\bm{c})}{\partial\bm{z}_{t}}.$ (2)
The Jacobian term necessitates backpropagation through the MMDiT, which is computationally expensive. However, this scenario is analogous to Score Distillation Sampling (SDS) (Poole et al., 2022), where research shows that omitting the Jacobian term yields an effective gradient for optimization with diffusion models. Thus, we adopt the same strategy for optimization.
3.3 Degradation-Suppression Guidance
Figure 4: Left: Robustness of FLUX. Right: Impacts of blurring different features in FLUX.
MSA loss effectively facilitates the insertion of reference objects. However, due to the inherent stochasticity of the denoising and optimization process, the results sometimes suffer from degraded visual quality, manifesting as oversaturated colors and reduced identity consistency (see Fig. 5). To address this, we introduce Degradation-Suppression Guidance (DSG), inspired by negative prompting (Schramowski et al., 2023), defined as:
$\bm{v}^{\mathrm{dsg}}_{t}=\bm{v}_{\bm{\theta}+\Delta\bm{\theta}}(\bm{z}_{t},t,\bm{c},\bm{z}^{\text{subj}})+\eta\bigl(\bm{v}_{\bm{\theta}+\Delta\bm{\theta}}(\bm{z}_{t},t,\bm{c},\bm{z}^{\text{subj}})-\bm{v}^{\mathrm{neg}}_{\bm{\theta}+\Delta\bm{\theta}}(\bm{z}_{t},t,\bm{c},\bm{z}^{\text{subj}})\bigr),$ (3)
where $\bm{v}^{\text{neg}}_{\bm{\theta}+\Delta\bm{\theta}}$ denotes a negative velocity prediction that guides the generation toward low-quality regions. A key challenge is the design of a meaningful negative velocity prediction $\bm{v}^{\text{neg}}_{\bm{\theta}+\Delta\bm{\theta}}$ within MMDiT-based architectures. In our experiments with FLUX, we observed that using nonsensical text prompts or explicit negative prompts fails to introduce degradation. The generated images remain high-fidelity (Fig. 4(a)), suggesting that text-based negative prompting is ineffective for FLUX.
In our setting, the ideal negative velocity $\bm{v}^{\text{neg}}_{\bm{\theta}+\Delta\bm{\theta}}$ should target directions that preserve the semantic content and spatial layout while lowering perceptual quality. To achieve this, we investigate whether we can manipulate FLUX’s internal representations to construct such a targeted degradation signal.
In FLUX, both multi-stream and single-stream blocks compute joint self-attention over concatenated text and image tokens as follows:
$\bm{h}=\text{softmax}\Big([\bm{Q}_{\text{txt}},\bm{Q}_{\text{img}}][\bm{K}_{\text{txt}},\bm{K}_{\text{img}}]^{\mathsf{T}}/\penalty 50\sqrt{d_{k}}\Big)\cdot[\bm{V}_{\text{txt}},\bm{V}_{\text{img}}],$ (4)
where $\bm{Q}_{\text{txt}},\bm{Q}_{\text{img}}]$ represents the concatenation of text and image queries, and similarly for keys and values. To identify an effective manipulation strategy, we systematically perturb different components in the attention mechanism (i.e., $\bm{Q}_{\text{txt}}$, $\bm{K}_{\text{txt}}$, $\bm{V}_{\text{txt}}$, $\bm{Q}_{\text{img}}$, $\bm{K}_{\text{img}}$ and $\bm{V}_{\text{img}}$) and observe their impact on generation quality.
As shown in Fig. 4(b), our findings are as follows:
  • 1.
    Blurring $\bm{Q}_{\text{txt}}$, $\bm{K}_{\text{txt}}$, or $\bm{V}_{\text{txt}}$ has negligible impact on semantic fidelity and visual quality.
  • 2.
    Blurring $\bm{V}_{\text{img}}$ severely disrupts the output distribution, leading to unintelligible images.
  • 3.
    Blurring $\bm{K}_{\text{img}}$ moderately impacts quality, while the image remains visually acceptable.
  • 4.
    Blurring $\bm{Q}_{\text{img}}$ yields pronounced degradations while preserving structural integrity, making it the most effective lever for constructing a negative velocity.
Based on these insights, we construct the negative velocity prediction $\bm{v}^{\text{neg}}_{\bm{\theta}+\Delta\bm{\theta}}$ in Eqn. 3 by blurring $\bm{Q}_{\text{img}}$ within FLUX (see Fig. 3(c)). Moreover, we show that blurring $\bm{Q}_{\text{img}}$ is mathematically equivalent to blurring the self-attention weights, whereas blurring $\bm{K}_{\text{img}}$ or $\bm{V}_{\text{img}}$ is not (see Appendix C for the proof). This equivalence is consistent with the fact that attenuating self-attention activations suppresses informative interactions and thus degrades image quality (Lu et al., 2024a).
Figure 5: Comparison of rectangular-mask blending and Adaptive Background Blending (ABB). Boundary regions (pink dashed boxes) are enlarged for clarity. Zoom in for details.
Algorithm 1 The Image Composition Process of SHINE.
1:Input: A background latent $\bm{z}^{\text{bg}}$, a inpainted latent $\bm{z}^{\text{init}}$, a subject latent $\bm{z}^{\text{subj}}$, a user mask $\bm{M}^{\text{user}}$.
2:Output: The composition latent $\bm{z}_{0}$}.
[…]
4:for $t={t_{1}},\dotsc,0$ do
5:  // Manifold-Steered Anchor (MSA) Optimization
6:  if $t>\tau$ then
7:   $\tilde{\bm{v}}_{t}\leftarrow\bm{v}_{\bm{\theta}}({\bm{z}}_{t},t,\bm{c})$
8:   for $j=1,\dotsc,k$ do
9:     $\bm{z}_{t}\leftarrow\bm{z}_{t}-\alpha\cdot\bm{M}^{\text{user}}\odot{\nabla_{\bm{z}_{t}}\left\|\bm{v}_{\bm{\theta}+\bm{\Delta\theta}}(\bm{z}_{t},t,\bm{c},\bm{z}^{\text{subj}})-{\tt sg}[\tilde{\bm{v}}_{t}]\right\|_{2}^{2}}$
10:   end for
11:  end if
12:  // Degradation-Suppression Guidance (DSG)
13:  $\bm{v}_{t},\bm{A}_{t}\leftarrow\bm{v}_{\bm{\theta}+\bm{\Delta\theta}}(\bm{z}_{t},t,\bm{c},\bm{z}^{\text{subj}})$
14:  ${\bm{v}}^{\text{dsg}}_{t}\leftarrow\bm{v}_{t}+\eta\left(\bm{v}_{t}-\bm{v}^{\text{neg}}_{\bm{\theta}+\bm{\Delta\theta}}(\bm{z}_{t},t,\bm{c},\bm{z}^{\text{subj}})\right)$
15:  $\bm{z}_{t-1}\leftarrow\bm{z}_{t}+(\sigma_{t-1}-\sigma_{t}){\bm{v}}^{\text{dsg}}_{t}$
16:  $\bm{z}^{\text{bg}}_{t-1}\leftarrow(1-\sigma_{t-1})\bm{z}^{\text{bg}}+\sigma_{t-1}\bm{\epsilon},~\text{where}~\bm{\epsilon}\sim\mathcal{N}\left(\mathbf{0},\mathbf{1}\right)$
17:  // Adaptive Background Blending (ABB)
18:  $\bm{M}^{\text{attn}}\leftarrow\text{MaxConnectedComponent}\big(\text{Dilate}(\mathbbm{1}(\bm{A}_{t}\geq\gamma))\big)$
19:  $\hat{\bm{M}}\leftarrow\mathbbm{1}{\{t>\tau\}}\;\bm{M}^{\text{attn}}+\mathbbm{1}{\{t\leq\tau\}}\;\bm{M}^{\text{user}}$
20:  $\bm{z}_{t-1}\leftarrow\hat{\bm{M}}\odot\bm{z}_{t-1}+\left(1-\hat{\bm{M}}\right)\odot\bm{z}^{\text{bg}}_{t-1}$
21:end for
22:return $\bm{z}_{0}$}
3.4 Adaptive Background Blending
Previous methods typically rely on the user-provided mask $\bm{M}^{\text{user}}$ to preserve the background during each denoising step, blending as $\bm{z}_{t}=\bm{M}^{\text{user}}\odot\bm{z}_{t}+(1-\bm{M}^{\text{user}})\odot\bm{z}^{\text{bg}}_{t}$, but this often introduces visible seams along mask boundaries (see the first row of Fig. 5).
To address this limitation, we propose Adaptive Background Blending (ABB), defined as
$\bm{z}^{\prime}_{t}=\hat{\bm{M}}\odot\bm{z}_{t}+\left(1-\hat{\bm{M}}\right)\odot\bm{z}^{\text{bg}}_{t},\;\;\;\;\;\;\;\;\;\;\;\hat{\bm{M}}=\mathbbm{1}{\{t>\tau\}}\;\mathcal{D}(\bm{M}^{\text{attn}})+\mathbbm{1}{\{t\leq\tau\}}\;\bm{M}^{\text{user}},$ (5)
where $\bm{M}^{\text{user}}$ is the user mask, while $\bm{M}^{\text{attn}}$ is derived by binarizing the cross-attention maps corresponding to subject tokens. These maps can be obtained by either averaging across layers or selecting the most informative layer via a lightweight analysis (details in Appendix D). The operator $\mathcal{D}(\cdot)$ performs dilation and extracts the largest connected component, ensuring robustness to noise.
Compared to $\bm{M}^{\text{user}}$, $\bm{M}^{\text{attn}}$ is more spatially precise, particularly for elongated or irregularly shaped objects that do not fully occupy a rectangular region. As illustrated in the second row of Fig. 5, our method produces smoother transitions by replacing the rigid user mask with the semantically guided mask. This refinement better preserves the surrounding scene, enabling seamless integration between generated content and the original background. However, applying this method throughout the denoising process may truncate object shadows or reflections. Through empirical evaluation, we find that leveraging $\bm{M}^{\text{attn}}$ during the initial denoising steps $(t>\tau)$ sufficiently mitigates visible seams along mask boundaries, ensuring high-fidelity scene coherence.
Figure 6: Qualitative comparison of our method with multiple baselines in challenging scenarios, drawn from our benchmark dataset. More qualitative comparisons are available in Appendix K.
4 Experiments
4.1 Experimental Setup
Benchmark. Current benchmarks primarily consist of background images with a fixed resolution of $512\times 512$ pixels. To assess performance across diverse, high-resolution, and complex scenarios, we introduce ComplexCompo, a benchmark built upon DreamEditBench (Li et al., 2023b). DreamEditBench includes 220 (subject, background, bounding box) pairs designed for $512\times 512$ resolution. In contrast, ComplexCompo features 300 composition pairs with varying resolutions, encompassing both landscape and portrait orientations, and incorporates challenging conditions such as low lighting, intense illumination, intricate shadows, and water reflections. The background images are sourced from OpenImage (Kuznetsova et al., 2020). Further details are provided in Appendix G.
Metrics. Previous methods primarily adapt CLIP-I (Radford et al., 2021) and DINOv2 (Oquab et al., 2023) to assess subject identity consistency. However, these features capture high-level semantic information that may not fully align with human perception of finer-grained attributes. Thus, we further incorporate instance retrieval features (IRF) from (Shao & Cui, 2022) and DreamSim (Fu et al., 2023), which better align with human judgments. An analysis of identity consistency metrics is provided in Appendix I. To evaluate overall image quality, we use ImageReward (IR) (Xu et al., 2023) and VisionReward (VR) (Xu et al., 2024), which are fine-grained, multi-dimensional reward models that more accurately reflect human preferences. For background consistency, LPIPS (Zhang et al., 2018) and SSIM (Wang et al., 2004) are employed.
Implementation Details. In our experiment, we used FLUX.1-dev, a 12B-parameter flow matching model, as the base model, combined with InstantCharacter (Tao et al., 2025) as the adapter (Additional results on SDXL, SD3.5, and PixArt are presented in Appendix E). Our approach also supports per-concept LoRA (Hu et al., 2021a), which requires test-time tuning (Ruiz et al., 2023b) but delivers superior identity consistency compared to an open-domain adapter, making it ideal for scenarios demanding precise identity preservation. The denoising process consists of 20 steps, with the inpainted image perturbed to timestep 15 and denoising initiated from that point. Additional details and hyperparameters are provided in Appendix J.
Table 1: Comparison of composition performance across two benchmarks. The best result in each column is highlighted in bold, while the second-best is underlined. Metrics shown in pink are those specifically trained to better align with human preferences. Abbreviations: IRF: Instance Retrieval Features; IR = ImageReward; VR = VisionReward.
Bench Method
Training
-Free
Base
Model
External
Model
Subject Identity Consistency Background Image Quality
CLIP-I $\uparrow$ DINOv2 $\uparrow$ IRF $\uparrow$ DreamSim $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ IR $\uparrow$ VR $\uparrow$
Dream-
Edit-
Bench
MADD (He et al., 2024) SD DINO 0.7118 0.6279 0.4333 0.5810 0.0604 0.8182 -0.2545 2.7011
ObjectStitch (Song et al., 2023) SD VIT 0.7567 0.6930 0.5525 0.5093 0.0190 0.8316 0.0791 3.2416
DreamCom (Lu et al., 2023c) SD LoRA 0.7414 0.6749 0.5597 0.5626 0.0200 0.8283 0.1873 3.5053
AnyDoor (Chen et al., 2024c) SD DINO 0.8183 0.7283 0.7714 0.3764 0.0251 0.8894 0.4511 3.3946
UniCombine (Wang et al., 2025) FLUX LoRA 0.8058 0.7332 0.7579 0.3984 0.0050 0.9397 0.4565 3.6108
PBE (Yang et al., 2023) SD - 0.7742 0.7040 0.5845 0.4985 0.0197 0.8287 0.2083 3.3482
TIGIC (Li et al., 2024b) SD - 0.7226 0.6718 0.4711 0.6108 0.0584 0.8153 -0.1332 2.9873
TALE (Pham et al., 2024) SD - 0.7329 0.6604 0.5007 0.6176 0.0392 0.8251 -0.1502 3.1349
TF-ICON (Lu et al., 2023d) SD - 0.7479 0.6865 0.5179 0.5441 0.0582 0.8111 0.0816 3.2823
DreamEdit (Li et al., 2023b) SD LoRA, VIT 0.7703 0.7151 0.6147 0.5047 0.0140 0.9775 0.1744 3.1775
EEdit (Yan et al., 2025) FLUX - 0.6998 0.6590 0.4438 0.6160 0.0039 0.9475 0.0216 3.3606
Ours-Adapter FLUX Adapter 0.8086 0.7415 0.7702 0.3730 0.0236 0.8959 0.5709 3.6234
Ours-LoRA FLUX LoRA 0.8125 0.7452 0.7900 0.3577 0.0271 0.8847 0.5906 3.6161
Complex-
Compo
MADD (He et al., 2024) SD DINO 0.6780 0.5993 0.3638 0.5979 0.0781 0.5658 -0.0088 2.6582
ObjectStitch (Song et al., 2023) SD VIT 0.7608 0.7077 0.5513 0.4717 0.0388 0.6357 0.2482 3.4411
DreamCom (Lu et al., 2023c) SD LoRA 0.648 0.5692 0.2788 0.8192 0.0389 0.6342 -0.0778 3.4409
AnyDoor (Chen et al., 2024c) SD DINO 0.7982 0.7052 0.7319 0.4493 0.0299 0.7262 0.3804 3.3787
UniCombine (Wang et al., 2025) FLUX LoRA 0.7361 0.6552 0.5380 0.5682 0.0237 0.7077 0.2470 3.5454
PBE (Yang et al., 2023) SD - 0.7537 0.6802 0.5189 0.5187 0.0397 0.6321 0.2139 3.4310
TIGIC (Li et al., 2024b) SD - 0.6913 0.6329 0.3848 0.6549 0.0929 0.6228 -0.131 2.8898
TALE (Pham et al., 2024) SD - 0.6816 0.6151 0.3799 0.6773 0.059 0.6334 0.0783 3.4498
TF-ICON (Lu et al., 2023d) SD - 0.6987 0.6435 0.4167 0.6030 0.0815 0.6216 0.1798 3.4323
DreamEdit (Li et al., 2023b) SD LoRA, VIT 0.7314 0.6722 0.5069 0.5670 0.0468 0.7201 0.1212 3.2531
EEdit (Yan et al., 2025) FLUX - 0.6713 0.6153 0.3797 0.6821 0.0226 0.7107 0.1433 3.5009
Ours-Adapter FLUX Adapter 0.7721 0.7107 0.6764 0.4294 0.0404 0.7789 0.4090 3.6020
Ours-LoRA FLUX LoRA 0.7999 0.7384 0.7659 0.3542 0.0430 0.7634 0.4246 3.5951
4.2 Experimental Results
We compare our method with two categories of baselines: (1) Training-based methods (6 in total): UniCombine (Wang et al., 2025), AnyDoor (Chen et al., 2024c), Paint by Example (PBE) (Yang et al., 2023), ObjectStitch (Song et al., 2023), MADD (He et al., 2024), and DreamCom (Lu et al., 2023c); (2) Training-free methods (5 in total): EEdit (Yan et al., 2025), TIGIC (Li et al., 2024b), DreamEdit (Li et al., 2023b), TF-ICON (Lu et al., 2023d), and TALE (Pham et al., 2024).
As shown in Tab. 1, both variants of our method surpass all baselines on DreamEditBench across human preference aligned metrics (i.e., DreamSim, IR, VR), which are the most critical indicators of quality. For background-related metrics, all methods achieve comparable results, with differences so small they are imperceptible to the human eye. On the more challenging ComplexCompo dataset, which includes non-square resolutions and intricate scenes, most methods experience a notable performance drop, yet our approach consistently remains the top performer. From Fig. 6, it is evident that while AnyDoor achieves high scores on many identity metrics, the model tends to copy and paste the subject into the scene, resulting in unnatural compositions and lower image quality scores. In contrast, our method excels at naturally composing objects in challenging conditions (e.g., low-light settings, water surfaces, and scenes with complex shadows). Appendix F provides user study.
4.3 Ablation Study
Figure 7: Qualitative ablation study comparing different variants of our framework.
Table 2: Ablation study examining the impact of key components on DreamEditBench.
Method
MSA
DSG
ABB
Subject Identity Consistency Background Image Quality
CLIP-I DINOv2 IRF DreamSim LPIPS SSIM IR VR
Config A 0.7328 0.6745 0.5754 0.5233 0.0166 0.9076 0.5577 3.5997
Config B 0.7814 0.7204 0.7414 0.3951 0.0172 0.9075 0.5455 3.5952
Config C 0.7528 0.6941 0.6533 0.4436 0.0178 0.9038 0.5633 3.6130
Config D 0.7421 0.6814 0.6158 0.5127 0.0210 0.9010 0.5595 3.6109
Config E 0.7481 0.6987 0.6647 0.4317 0.0218 0.8971 0.5850 3.6277
Config F 0.8084 0.7429 0.7609 0.3756 0.0231 0.8991 0.5459 3.6023
Config G 0.8077 0.7375 0.7589 0.3762 0.0182 0.9037 0.5745 3.6191
Ours-Adapter 0.8086 0.7415 0.7702 0.3730 0.0236 0.8959 0.5709 3.6232
We validate our design choices through ablation (see Tab. 2 and Fig. 7). The results highlight three key insights. First, MSA loss notably improves subject identity consistency. Second, DSG boosts IR and VR scores by steering denoising away from low-quality regions. Finally, ABB effectively suppresses visible seams along mask boundaries. While this improvement is readily apparent in visual comparisons (Figs. 5, 7), it is less well captured by quantitative metrics, since LPIPS and SSIM primarily assess structural similarity rather than perceptual smoothness.
5 Conclusion
We introduced SHINE, a training-free framework for seamless and high-fidelity image composition with pretrained T2I models. SHINE integrates Manifold-Steered Anchor Loss, Degradation-Suppression Guidance, and Adaptive Background Blending to ensure precise subject placement and artifact-free synthesis across diverse resolutions and lighting conditions. To enable rigorous evaluation, we proposed ComplexCompo, a benchmark for challenging composition scenarios. SHINE achieves state-of-the-art results on both ComplexCompo and DreamEditBench.
Figure 8: Composition inherit erroneous colors if the inpainting prompt specifies an incorrect color.
Limitations. Our method reliably converges to the correct subject identity through MSA optimization even when the inpainted subject’s appearance deviates substantially from the reference (see Fig. 8(a)). However, when the inpainting prompt specifies an incorrect color, the final inpainted result tends to inherit and preserve this erroneous color (see Fig. 8(b)). On the other hand, the similarity between the inserted object and the user-provided object depends on the quality of the customization adapter used. As shown in Tab. 1, because LoRA performs test-time tuning for individual concepts, it generates subjects that are more similar to the target than those produced by pretrained open-domain customization adapters, resulting in higher subject identity consistency metrics in the composition. While current customization adapters already perform well, the potential of our method will continue to improve as advancements are made in the field of open-domain customization adapters.
Appendix
Appendix A Related Work
A.1 Image Composition
Image composition involves integrating specific objects and scenarios from user-provided photos, often guided by text prompts. Traditionally, this process is divided into sub-tasks (Niu et al., 2021), including object placement (Azadi et al., 2020; Chen & Kae, 2019; Lin et al., 2018; Tripathi et al., 2019; Zhang et al., 2020a), image blending (Wu et al., 2019; Zhang et al., 2020b), image harmonization (Cao et al., 2023; Lu et al., 2023b; Zhang et al., 2020b; Cong et al., 2020; Cun & Pun, 2020; Jiang et al., 2021; Xue et al., 2022), and shadow generation (Hong et al., 2022; Liu et al., 2020; Sheng et al., 2021; Zhang et al., 2019), each typically handled by distinct models and pipelines. However, with the rise of diffusion-based generative models, recent approaches have shifted toward unified frameworks that address all sub-tasks simultaneously. These methods are broadly classified into training-based and training-free approaches.
Training-based methods fine-tune foundational models using datasets tailored for image composition. Early methods like Paint by Example (Yang et al., 2022) and ObjectStitch (Song et al., 2022) use CLIP to encode subject features, ensuring high semantic similarity between inserted objects and reference images. These approaches use image augmentation to create training datasets, enabling effective training. GLIGEN (Li et al., 2023c) incorporates grounding information into new trainable layers of a pre-trained diffusion model via a gated mechanism. ControlCom (Zhang et al., 2023a) integrates 2-dim indicator vector to improve controllability. DreamCom (Lu et al., 2023c) and MureObjectStitch (Chen et al., 2024a) fine-tune models with small sets of reference images to preserve subject identity. AnyDoor (Chen et al., 2024c), IMPRINT (Song et al., 2024b), and E-MD3C (Pham et al., 2025) leverage DINOv2 to enhance identity fidelity and control over shape and pose, drawing supervision from video data. MimicBrush (Chen et al., 2024b) similarly uses video-derived supervision for imitative editing. In contrast, MADD (He et al., 2024), ObjectMate (Winter et al., 2024), and OmniPaint (Yu et al., 2025b) employ image inpainting models to generate higher-quality training datasets compared to those based on image or video augmentation. Multitwine (Tarrés et al., 2025) enables the integration of multiple objects, capturing interactions from simple positional relationships to complex actions requiring reposing. DreamFuse (Huang et al., 2025) uses a Positional Affine mechanism to embed foreground size and position into the background, fostering effective foreground-background interaction through shared attention. Insert Anything (Song et al., 2025) and UniCombine (Wang et al., 2025) introduces a FLUX-based, multi-conditional generative framework that handles diverse condition combinations. However, these methods often bias the generative priors of base models toward curated datasets, resulting in unnatural compositions, such as implausible object-environment interactions (e.g., missing or unrealistic shadows and reflections). This stems from the absence of a large-scale, high-quality, multi-resolution, real-world triplet dataset comprising an object, a scene, and the object seamlessly integrated into the scene, which is expensive to produce.
Training-free methods, on the other hand, modify the inference process of pre-trained models to achieve composition without additional training. Early approaches like TF-ICON (Lu et al., 2023d) leverage accurate image inversion to lay the groundwork for composition, achieved through composite self-attention map injection. TALE (Pham et al., 2024) and PrimeComposer (Wang et al., 2024c) build on TF-ICON to enhance identity preservation and background-object style adaptation. TIGIC (Li et al., 2024b) focuses on preserving non-target areas during composition. Thinking Outside the BBox (Tarrés et al., 2024) enables unconstrained image compositing, unbound by input masks. FreeCompose (Chen et al., 2024e) employs a pipeline of object removal, image harmonization, and semantic composition. DreamEdit (Li et al., 2023b), UniCanvas (Jin et al., 2025), and Magic Insert (Ruiz et al., 2024) use test-time tuning to fine-tune models during inference. Add-it (Tewel et al., 2024) enables text-guided object insertion on FLUX, where users describe objects via text prompts instead of reference images. EEdit (Yan et al., 2025) recently improves TF-ICON on FLUX, introducing step-skipping to reduce time costs and spatial locality caching to minimize redundancy. However, training-free methods rely on precise image inversion and fragile attention surgery, which can lock inserted objects into the exact pose of the reference image, leading to awkward or contextually inappropriate orientations. Attention manipulation often causes instability and hyperparameter sensitivity, as feature or attention map injection does not always preserve subject identity. This creates a trade-off: stronger injection preserves identity but results in unnatural poses, while lighter injection yields more natural poses but compromises identity.
A.2 General Image Editing
Recent and significant advancements in text-to-image generative models have enhanced numerous applications (Zhou et al., 2025c; a; b; 2023; 2024a; 2024b; 2024c; Zhao et al., 2024a; d; c; 2025b; 2025a; b; Xie et al., 2024; Wang et al., 2023b; Wang & Chen, 2025a; b; Lu et al., 2024b; Li et al., 2025; Gao et al., 2024; 2025; Zhu et al., 2024; Yu et al., 2025a). Instruction-based image editing has evolved rapidly. Early systems relied on modular, two-stage pipelines: a multimodal language model first produced textual prompts, spatial guidance, or synthetic instruction–image pairs, and a separate diffusion model then executed the edit—as in InstructEdit (Wang et al., 2023a), InstructPix2Pix (Brooks et al., 2022), MagicBrush (Zhang et al., 2023b), and BrushEdit (Li et al., 2024c). Recent work has shifted toward tightly integrated, instruction-centric architectures. Models such as SmartEdit (Huang et al., 2024), X2I (Ma et al., 2025), RPG (Yang et al., 2024), AnyEdit (Yu et al., 2024), and UltraEdit (Zhao et al., 2024e) embed routing, task-aware objectives, and fine-grained controls directly into the network, yielding higher fidelity and more precise manipulation.
Unified generation-and-editing frameworks (e.g., OmniGen (Xiao et al., 2024), ACE (Han et al., 2024), ACE++ (Mao et al., 2025), Lumina-OmniLV (Pu et al., 2025), Qwen2VL-Flux (Lu, 2024), DreamEngine (Chen et al., 2025), MetaQueries (Pan et al., 2025), Hidream-E1 (HiDream-ai, 2025)) treat editing as one capability of an end-to-end vision-language model, often fusing language embeddings with diffusion latents to provide context-aware, pixel-level control. Efficiency has advanced in parallel: ICEdit (Zhang et al., 2025) couples LoRA with mixture-of-experts tuning and optimized noise initialization, while SuperEdit (Ming et al., 2025) relies on higher-quality data and contrastive supervision to sustain performance at lower cost. Looking ahead, large foundation models such as Gemini (Gemini2.5, 2025) and GPT-5 (OpenAI, 2025) already show strong visual reasoning and coherent, instruction-guided image generation. Yet, despite extensive multimodal pre-training, they still fall short on image composition: object placement remains hard to control, lighting is often inconsistent, and subjects can drift in identity.
A.3 Subject-Driven Generation
Extensive research has explored subject-driven image generation, in which the output must not only portray the contexts described by the text prompt but also faithfully include the specific subject supplied by reference images. Methods in this area are divided into two categories—test-time fine-tuning customization and zero-shot customization—according to whether extra training is needed for each new subject. Our framework accommodates both categories, so we provide two corresponding variants in the main paper.
Test-time fine-tuning methods (Gal et al., 2022a; Ruiz et al., 2023a) adapt a pre-trained T2I model to a small set of reference images (typically 3 to 5 images). Although this step adds computational cost and latency, it offers the greatest flexibility for diverse customization requirements. Such methods are commonly grouped into three subclasses: data regularization, weight regularization, and loss regularization. In the data-regularization family, DreamBooth (Ruiz et al., 2023a) limits overfitting by generating superclass images with the base T2I model and training on both reference and regularization images; Custom Diffusion (Kumari et al., 2023) improves regularization quality by retrieving real images; and Specialist Diffusion (Lu et al., 2023a) applies extensive data augmentation. Weight-regularization approaches (Gal et al., 2022b; Hu et al., 2021b; Han et al., 2023; Qiu et al., 2023) confine updates to carefully chosen parameters, such as the subject-specific text embeddings or the singular values of weight matrices. Loss-regularization approaches, including Specialist Diffusion (Lu et al., 2023a), MagiCapture (Hyung et al., 2023), and FaceChain-SuDe (Qiao et al., 2024), introduce objective terms that respectively maximize CLIP-space similarity between generated and reference images, disentangle identity and style via masked facial reconstruction, or encourage correct superclass classification.
Zero-shot image customization methods avoid subject-specific fine-tuning at inference time but rely on extensive pre-training. For general subject customization, InstantBooth (Shi et al., 2024a) adds a visual encoder that captures coarse-to-fine image features from the references; BLIP-Diffusion (Li et al., 2024a) fine-tunes BLIP-2 (Li et al., 2023a) to extract multimodal subject representations; ELITE (Wei et al., 2023) maps reference images into hierarchical textual tokens through global and local networks; and Song et al. (Song et al., 2024a) enhance textual control by removing the projection of visual embeddings onto textual embeddings. For facial customization, InstantID (Wang et al., 2024b) isolates facial regions from reference images to extract appearance and structural cues. For style customization, InstantStyle (Wang et al., 2024a) identifies style-controlling layers and injects IP-Adapter features (Ye et al., 2023a) into those layers to achieve style transfer. InstantCharacter (Tao et al., 2025), IP-Adapter (Ye et al., 2023b), and PuLID (Guo et al., 2024) have each released versions compatible with the FLUX model.
Appendix B Visualizing the Impact of Adaptive Background Blending
Although our loss function aims to find a new latent within the manifold learned by the adapter, encouraging the adapter-augmented T2I model’s predictions to closely match those of the base model on the original noisy latent, this early-stage optimization primarily preserves structural elements rather than fine details. Consequently, discrepancies in fine-grained features often arise between the masked and unmasked regions. As illustrated in Fig. 9, we compare the composite images generated using our Adaptive Background Blending (ABB) method with those produced via direct background blending using the rectangular user mask. For clarity, we enlarge the boundary areas of each image (highlighted in pink dashed boxes) to better reveal differences.
Figure 9: Comparison of composites from our Adaptive Background Blending (ABB) method and direct blending with a rectangular mask. Boundary regions within pink dashed boxes are enlarged for clarity. Please zoom in to see details.
Appendix C Equivalence of Query Blurring and Attention Weight Blurring
[…]
C.1 Blurring the Query Matrix
Blurring the self-attention weights $\bm{QK}^{\mathsf{T}}$ with a 2D Gaussian filter $\bm{G}$ can be expressed as:
$\bm{G}\circledast(\bm{QK}^{\mathsf{T}}),$ (6)
[…]
$\bm{G}\circledast(\bm{QK}^{\mathsf{T}})=\bm{B}(\bm{QK}^{\mathsf{T}}).$ (7)
Using the properties of matrix multiplication, we can rewrite this as:
$\bm{B}(\bm{QK}^{\mathsf{T}})=(\bm{BQ})\bm{K}^{\mathsf{T}}.$ (8)
Since the convolution operation is linear, applying the Gaussian filter $\bm{G}$ to the rows of $\bm{Q}$ yields:
$\bm{BQ}=\bm{G}\circledast\bm{Q}.$ (9)
Thus, we obtain:
$\bm{G}\circledast(\bm{QK}^{\mathsf{T}})=(\bm{G}\circledast\bm{Q})\bm{K}^{\mathsf{T}}.$ (10)
This establishes that blurring the query matrix $\bm{Q}$ with $\bm{G}$ is mathematically equivalent to applying the same blur to the self-attention weights $\bm{QK}^{\mathsf{T}}$. This equivalence suggests that query blurring can be used as a computationally efficient proxy for smoothing attention weights, potentially reducing the need for direct manipulation of the attention matrix.
C.2 Blurring the Key and Value Matrices
In contrast, applying the Gaussian blur to the key matrix $K$ does not yield a similar equivalence. Consider the convolution applied to $K$. The resulting attention weights become:
$\bm{Q}(\bm{G}\circledast\bm{K})^{\mathsf{T}}=\bm{Q}(\bm{BK})^{\mathsf{T}}=\bm{Q}\bm{K}^{\mathsf{T}}\bm{B}^{\mathsf{T}}.$ (11)
Since $\bm{B}^{\mathsf{T}}\neq\bm{B}$ for a general Toeplitz matrix derived from a Gaussian filter, we have:
$\bm{QK}^{\mathsf{T}}\bm{B}^{\mathsf{T}}\neq\bm{B}(\bm{QK}^{\mathsf{T}}).$ (12)
Thus, blurring the key matrix $\bm{K}$ does not produce an equivalent effect to blurring the attention weights $\bm{QK}^{\mathsf{T}}$. A similar argument applies to the value matrix $\bm{V}$, as the output of self-attention, $(\bm{QK}^{\mathsf{T}})\bm{V}$, involves $\bm{V}$ in a post-multiplication step, and convolution on $\bm{V}$ does not commute with the attention weight computation in the same manner.
Appendix D Evaluating Cross-Attention Map Accuracy via IoU
To identify the most accurate cross-attention maps that reflect the location of the generated object, we first create 100 prompts containing a main subject (e.g.,“a dog is sleeping on a couch”) using GPT-5. These prompts are then used to generate 100 images with FLUX.1-dev, employing 20 denoising steps. Cross-attention maps are extracted from 19 multi-stream (or double-stream) blocks and 38 single-stream blocks across all denoising steps. The maps are averaged over the 20 steps and subsequently normalized and binarized, resulting in a total of 57 binary masks.
To determine which of these 57 masks is the most accurate, we compute the Intersection over Union (IoU) between each mask and a ground-truth mask obtained by segmenting the final generated images using SAM (Ravi et al., 2025). The IoU for each block is averaged over the 100 generated images. The results are presented in Fig. 10, showing that the cross-attention maps from the last multi-stream (or double-stream) block achieve the highest segmentation accuracy. A visualization of the cross-attention maps from all blocks is provided in Fig. 11.
Figure 10: The IoU is calculated between the mask produced from each block and the ground-truth mask, which is obtained by segmenting the final generated images using SAM. The IoU for each block is averaged over 100 images.
Figure 11: Visualization of cross-attention maps from different MMDiT blocks of FLUX.1-dev.
Appendix E Experiments with SDXL, SD3.5, and PixArt
Our framework introduces a novel, model-agnostic approach for enhancing generative architectures, anchored by three synergistic components: Manifold-Steered Anchor (MSA) loss, Degradation-Suppression Guidance (DSG), and Adaptive Background Blending (ABB). These components are meticulously designed to leverage ubiquitous features of modern generative models, ensuring seamless integration without requiring architectural modifications. Specifically, MSA utilizes either LoRA-based personalization or a pretrained personalization adapter, DSG capitalizes on widely available self-attention maps, and ABB harnesses text–image cross-attention maps, a staple in most text-to-image pipelines. This design ensures broad applicability across diverse generative models.
In the main text, we showcase the effectiveness of our approach on Flux, a leading open-source model. To establish its generalizability, we conduct experiments on SDXL (Podell et al., 2024), SD3.5 (Esser et al., 2024), and PixArt (Chen et al., 2023) across DreamEditBench and ComplexCompo. The results are presented in Tab. 3. On DreamEditBench, our LoRA-based variant with Flux achieves state-of-the-art performance in subject identity preservation, while delivering superior image quality. Similarly, on ComplexCompo, the Flux-LoRA configuration excels in identity consistency and image fidelity. Notably, the framework’s benefits extend beyond a single model family: both SDXL and PixArt-$\Sigma$ exhibit substantial performance gains, affirming the approach’s generality and adaptability across diverse generative architectures.
Table 3: Comparison of compositional performance across two benchmarks with different base models. The best result in each column is highlighted in bold, while the second-best is underlined. Metrics shown in pink are those specifically trained to better align with human preferences. Abbreviations: IRF: Instance Retrieval Features; IR = ImageReward; VR = VisionReward.
Bench Method
Base
Model
Subject Identity Consistency Background Image Quality
CLIP-I $\uparrow$ DINOv2 $\uparrow$ IRF $\uparrow$ DreamSim $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ IR $\uparrow$ VR $\uparrow$
DreamEdit-
Bench
Ours-Adapter SDXL 0.7944 0.7334 0.7659 0.3761 0.0238 0.8922 0.5621 3.6158
Ours-Adapter SD3.5 0.8054 0.7407 0.7699 0.3745 0.0234 0.8937 0.5701 3.6187
Ours-LoRA PixArt-$\Sigma$ 0.8098 0.7445 0.7798 0.3612 0.0251 0.8875 0.5842 3.6198
Ours-Adapter FLUX 0.8086 0.7415 0.7702 0.3730 0.0236 0.8959 0.5709 3.6234
Ours-LoRA FLUX 0.8125 0.7452 0.7900 0.3577 0.0271 0.8847 0.5906 3.6161
Complex-
Compo
Ours-Adapter SDXL 0.7657 0.7084 0.6862 0.4457 0.0457 0.7612 0.3894 3.5987
Ours-Adapter SD3.5 0.7701 0.7091 0.6977 0.4173 0.0401 0.7784 0.4091 3.6021
Ours-LoRA PixArt-$\Sigma$ 0.7924 0.7287 0.7311 0.3603 0.0424 0.7698 0.4277 3.5988
Ours-Adapter FLUX 0.7721 0.7107 0.6764 0.4294 0.0404 0.7789 0.4090 3.6020
Ours-LoRA FLUX 0.7999 0.7384 0.7659 0.3542 0.0430 0.7634 0.4246 3.5951
Appendix F User Study
We conduct a user study involving 50 participants. Each participant was asked to complete 50 ranking tasks. In each task, they were shown 13 composition results generated by different methods, along with a reference subject image.
To ensure a balanced evaluation, 25 of the tasks were randomly sampled from DreamEditBench and the remaining 25 from ComplexCompo. Participants were asked to rank the results based on two key criteria: (1) subject identity consistency and (2) composition realism. A lower rank (e.g., 1st) indicates a better composition result, while a higher rank (e.g., 13th) reflects a less favorable outcome.
We summarize the average ranking scores for each method in Tab. 4. Our method received the most favorable rankings from the majority of participants, demonstrating its effectiveness in producing high-quality compositions.
Table 4: Average ranking scores from the user study on image composition methods. Lower is better.
Method Training-Free Average Ranking (Lower is Better)
MADD 12.44
ObjectStitch 11.80
DreamCom 6.66
AnyDoor 4.12
UniCombine 2.94
PBE 4.94
TIGIC 9.74
TALE 9.06
TF-ICON 8.36
DreamEdit 6.36
EEdit 10.76
Ours-Adapter 2.30
Ours-LoRA 1.52
Appendix G Benchmark Details
Our benchmark consists of 300 triplets, each comprising a subject image, a background image, and a bounding box. The subject images are identical to those used in DreamEditBench (Li et al., 2023b; Ruiz et al., 2023a), while the background images are sampled from the OpenImages dataset (Kuznetsova et al., 2020). These backgrounds exhibit a variety of aspect ratios and resolutions, including landscape and portrait formats, such as $768\times 1088$, $768\times 1072$, $768\times 1024$, $768\times 1152$, $024\times 768, 11$$152\times 768$, $1200\times 768$, $848\times 768$, and $1360\times 768$. The bounding boxes are manually designed to ensure that the size and placement of the inserted subjects are contextually appropriate and visually plausible. The benchmark will be released publicly.
Appendix H Prompts for Proprietary Foundation Models
To perform image composition with proprietary, general-purpose multimodal foundation models (e.g., GPT-5 (OpenAI, 2025), Gemini 2.5 Pro (Gemini2.5, 2025), SeedEdit/Doubao (Shi et al., 2024b), and Grok 4 (gro, 2025)), we upload three images: (1) Subject image; (2) Background image; and (3) Mask image defining the insertion region.
We then issue a prompt of the following form (with the resolution and coordinates adjusted for each case):
Please insert the object from the first uploaded image into the second image. The target region for insertion is defined by the mask in the third image. For reference, the resolution of the second image is $1152\times 768$, and the bounding box for placement is specified by the top-left and bottom-right coordinates: $(x_{1}=550,y_{1}=600,x_{2}=700,y_{2}=750)$. The inserted object should retain the same identity and appearance as in the first image. The final composite should appear realistic, natural, and physically plausible.
Appendix I Subject Identity Metrics Analysis
In our experiments we found that widely-used subject-identity metrics such as CLIP-I (Radford et al., 2021) and DINOv2 (Oquab et al., 2023) correlate poorly with human preferences. Because they focus almost exclusively on semantic similarity, they ignore appearance changes introduced by lighting, shadows, reflections, and surrounding context. Fig. 12(b) presents several image pairs produced by AnyDoor (left) and by our method (right); the corresponding CLIP-I $(\uparrow)$ (Radford et al., 2021), DINOv2 $(\uparrow)$ (Oquab et al., 2023), IRF $(\uparrow)$ (Shao & Cui, 2022), and DreamSim $(\downarrow)$ (Fu et al., 2023) scores are shown beneath each image, with the better score highlighted in red. Although the AnyDoor results are visibly less realistic and less consistent, they nevertheless receive higher CLIP-I and DINOv2 scores, and in most cases higher IRF scores, demonstrating that these measures do not faithfully capture compositional quality.
A reliable metric should recognise the same object whether it is underwater (see Fig. 12(b)[(8), (17), (20)]), in shadow (see Fig. 12(b)[(9), (13), (19)]), partially occluded (see Fig. 12(b)[(2)]), or situated in a low-light or back-lit scene (see Fig. 12(b)[(5), (14), (15), (16)]). Among the metrics we evaluated, only DreamSim, which was designed to align more closely with human perception, consistently exhibits this desired behaviour.
Figure 12: Comparison of Subject Identity Metrics. (a) Reference subject images used for metric calculations. (b) Image pairs generated by AnyDoor (left) and our method (right), with corresponding CLIP-I $(\uparrow)$, DINOv2 $(\uparrow)$, IRF $(\uparrow)$, and DreamSim $(\downarrow)$ scores displayed below each image; the better score is highlighted in red. Despite AnyDoor’s results appearing less realistic and consistent, they often achieve higher CLIP-I, DINOv2, and IRF scores, indicating that these metrics may not reliably reflect compositional quality. In contrast, DreamSim provides a more reliable assessment.
Appendix J Implementation Details
The hyperparameters used in our frameworks are summarized in Tab. 5. Each baseline is implemented according to the configuration settings recommended in its original publication. The repositories utilized for each baseline are listed below:
Table 5: Hyperparameters of Our Frameworks. Bin Thresh = Binarization Threshold; #iter = Number of Iterations.
Variant Denoising Setup Manifold-Steered Anchor Loss Degradation-Suppression Guidance Adaptive Background Blending
Total steps
$(T-1)\to 0$
Start
step $t_{1}$
Step Range
$t_{1}\to\tau$
Learning Rate
$\alpha$
#iters
$k$
Step Range
$t_{1}\to 0$
Scale
$\eta$
Blur
$\sigma$
Step Range
$(t_{1}-1)\to 1$
Bin
Thresh
Dilation
Kernel Size
Ours-Adapter $19\to 0$ 14 $14\to 12$ 500, 750, 1000 10 $14\to 0$ 0.5 10 $13\to 1$ 0.2 3
Ours-LoRA $19\to 0$ 13 $13\to 12$ 50, 300 2 $13\to 0$ 0.7 10 $12\to 1$ 0.4 3
Appendix K Additional Qualitative Results
We offer more qualitative assessment results, including visualizations of all baselines, presented in Figs. 13 to 18.
Appendix L LLM Usage Statement
We used large language models for text polishing and grammar correction during manuscript preparation. No LLMs were involved in the design of the method, experiments, or analysis. All content has been carefully verified and validated by the authors.
Figure 13: (Part 1 of 2) Qualitative comparison of our method against baselines in challenging scenarios.
Figure 14: (Part 2 of 2) Qualitative comparison of our method against baselines in challenging scenarios.
Figure 15: (Part 1 of 2) Qualitative comparison of our method against baselines in challenging scenarios.
Figure 16: (Part 2 of 2) Qualitative comparison of our method against baselines in challenging scenarios.
Figure 17: (Part 1 of 2) Qualitative comparison of our method against baselines in challenging scenarios.
Figure 18: (Part 2 of 2) Qualitative comparison of our method against baselines in challenging scenarios.

### v4

@@
-Does FLUX Already Know How to Perform Physically Plausible Image Composition?
+arXiv:2509.21278v4[cs.CV] 28 Feb 2026
 Abstract
@@
 Experiments on ComplexCompo and DreamEditBench show state-of-the-art performance on standard metrics (e.g., DINOv2) and human-aligned scores (e.g., DreamSim, ImageReward, VisionReward).
-Code and benchmark will be publicly available upon publication.
+Code is available at https://github.com/ZhumingLian/SHINE.
 Figure 1: Showcase of our training-free image composition method, SHINE.
@@
 Yet building large-scale, high-quality, multi-resolution triplet datasets (object, scene, composite) is prohibitively costly.
-As a result, most composition models are fine-tuned from base models (e.g., FLUX.1-dev (Black Forest Labs, 2024a), FLUX.1-Fill (Black Forest Labs, 2024c), SDXL (Podell et al., 2024)) on synthetic data generated via inpainting or augmentations (Chen et al., 2024c; Yang et al., 2023; Song et al., 2023; Wang et al., 2025; He et al., 2024).
+As a result, most composition models are fine-tuned from base models (e.g., FLUX.1-dev (Black Forest Labs, 2024a), FLUX.1-Fill (Black Forest Labs, 2024c), SDXL (Podell et al., 2024)) on synthetic data generated via inpainting or augmentations (Chen et al., 2024c; Yang et al., 2023; Song et al., 2023; Wang et al., 2025a; He et al., 2024).
 These models, however, face two main limitations (see Fig.
@@
 Notably, such issues are absent in the base models, implying that the underlying physical priors are present but are not effectively exploited by fine-tuned variants.
-The degradation largely stems from low-quality synthetic datasets, which inherit flaws from inpainting models that often mis-handle shadows and reflections, producing implausible edits, hallucinated content, or incomplete object removal (Yu et al., 2025b; Winter et al., 2024).
+The degradation largely stems from low-quality synthetic datasets, which inherit flaws from inpainting models that often mis-handle shadows and reflections, producing implausible edits, hallucinated content, or incomplete object removal (Yu et al., 2025c; Winter et al., 2025).
 There have been prior training-free attempts to exploit the priors of text-to-image (T2I) models for advancing image composition, but they fall short for two main reasons.
 (i) Inversion bottlenecks.
-Most methods (Lu et al., 2023d; Pham et al., 2024; Yan et al., 2025; Li et al., 2024b) depend on accurate image inversion (Song et al., 2020; Lu et al., 2022; Mokady et al., 2023).
+Most methods (Lu et al., 2023d; Pham et al., 2024; Yan et al., 2025; Li et al., 2024b) depend on accurate image inversion (Song et al., 2021; Lu et al., 2022; Mokady et al., 2023).
 In practice, inversion constrains the inserted object to the pose of its reference image, often resulting in contextually inappropriate orientations.
@@
 1).
-SHINE comprises three innovations: (i) Manifold-Steered Anchor (MSA) loss, which leverages pretrained open-domain customization adapters (e.g., IP-Adapter (Ye et al., 2023b)) to steer noisy latents toward faithfully representing the reference subject while preserving the structural integrity of the background.
+SHINE comprises three innovations: (i) Manifold-Steered Anchor (MSA) loss, which leverages pretrained open-domain customization adapters (e.g., IP-Adapter (Ye et al., 2023)) to steer noisy latents toward faithfully representing the reference subject while preserving the structural integrity of the background.
 (ii) Degradation-Suppression Guidance (DSG) that steers sampling away from low-quality distributions.
@@
 To evaluate performance across diverse, high-resolution, and demanding scenarios, we introduce ComplexCompo, a benchmark that includes varied resolutions, both landscape and portrait orientations, and complex conditions such as low lighting, intense illumination, intricate shadows, and water reflections.
-Extensive experiments on ComplexCompo and DreamEditBench (Li et al., 2023b) demonstrate that SHINE achieves state-of-the-art (SOTA) performance, surpassing baselines on standard metrics (e.g., DINOv2 (Oquab et al., 2023)) and human-aligned metrics (e.g., DreamSim (Fu et al., 2023), ImageReward (Xu et al., 2023), VisionReward (Xu et al., 2024)).
+Extensive experiments on ComplexCompo and DreamEditBench (Li et al., 2023b) demonstrate that SHINE achieves state-of-the-art (SOTA) performance, surpassing baselines on standard metrics (e.g., DINOv2 (Oquab et al., 2024)) and human-aligned metrics (e.g., DreamSim (Fu et al., 2023), ImageReward (Xu et al., 2023), VisionReward (Xu et al., 2024)).
 Figure 2: Image composition from advanced multimodal models under three challenging conditions: backlighting, shadows, and water surfaces. Refer to Appendix H for prompt details.
@@
 Diffusion models have shifted the field toward unified frameworks, either training-based or training-free.
-Training-based approaches fine-tune diffusion models with curated datasets, adding grounding layers, controllability signals, or identity-preserving supervision from image or video sets (Wang et al., 2025; Chen et al., 2024c; Yang et al., 2023; Song et al., 2023; Lu et al., 2023c).
+Training-based approaches fine-tune diffusion models with curated datasets, adding grounding layers, controllability signals, or identity-preserving supervision from image or video sets (Wang et al., 2025a; Chen et al., 2024c; Yang et al., 2023; Song et al., 2023; Lu et al., 2023c).
 However, they often bias model priors and struggle with complex lighting due to the lack of large-scale real-world triplets.
@@
 These mild assumptions enable seamless integration into existing pipelines without architectural changes.
-We present main results with FLUX, while additional experiments on SDXL (Podell et al., 2024), SD3.5 (Esser et al., 2024), and PixArt (Chen et al., 2023) are provided in Appendix E.
+We present main results with FLUX, while additional experiments on SDXL (Podell et al., 2024), SD3.5 (Esser et al., 2024), and PixArt (Chen et al., ) are provided in Appendix E.
 The complete algorithm is shown in Algorithm 1.
@@
 As illustrated in Fig.
-3(a), we use a vision-language model (VLM) (Xue et al., 2024; Chen et al., 2024d; Liu et al., 2024) to caption the subject image and leverage this caption, along with an image inpainting model (Li et al., 2024c; Ju et al., 2024; Zhuang et al., 2024; Black Forest Labs, 2024b), to generate the inpainted background image, denoted as $\bm{x}^{\text{init}}$.
+3(a), we use a vision-language model (VLM) (Xue et al., 2024; Chen et al., 2024d; Liu et al., 2024) to caption the subject image and leverage this caption, along with an image inpainting model (Li et al., 2024c; Ju et al., 2024; Zhuang et al., 2024; Black Forest Labs, 2024b), to generate the image to which the subject is attached, denoted as $\bm{x}^{\text{init}}$.
 The noisy latent is encoded in the VAE space as $\bm{z}^{\text{init}}$ and perturbed to timestep $t\leq T$ via one-step forward diffusion, following the flow matching formulation: $\bm{z}_{t}=(1-\sigma_{t})\bm{z}^{\text{init}}+\sigma_{t}\bm{\epsilon}$, where $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{1})$.
@@
 3.1) during the denoising process, steering it toward a reference subject while preserving the structural integrity of the original image.
-The key intuition is to leverage the prior knowledge embedded in pretrained open-domain customization adapters (or alternatively, personalized LoRAs) such as IP-Adapter (Ye et al., 2023b), PuLID (Guo et al., 2024), and InstantCharacter (Tao et al., 2025), to intervene directly in the diffusion trajectory.
+The key intuition is to leverage the prior knowledge embedded in pretrained open-domain customization adapters (or alternatively, personalized LoRAs) such as IP-Adapter (Ye et al., 2023), PuLID (Guo et al., 2024), and InstantCharacter (Tao et al., 2025), to intervene directly in the diffusion trajectory.
 Specifically, the MSA loss is defined as:
@@
 The Jacobian term necessitates backpropagation through the MMDiT, which is computationally expensive.
-However, this scenario is analogous to Score Distillation Sampling (SDS) (Poole et al., 2022), where research shows that omitting the Jacobian term yields an effective gradient for optimization with diffusion models.
+However, this scenario is analogous to Score Distillation Sampling (SDS) (Poole et al., ), where research shows that omitting the Jacobian term yields an effective gradient for optimization with diffusion models.
 Thus, we adopt the same strategy for optimization.
@@
 Algorithm 1 The Image Composition Process of SHINE.
-1:Input: A background latent $\bm{z}^{\text{bg}}$, a inpainted latent $\bm{z}^{\text{init}}$, a subject latent $\bm{z}^{\text{subj}}$, a user mask $\bm{M}^{\text{user}}$.
+1:Input: A background latent $\bm{z}^{\text{bg}}=\text{VAE}(\bm{x}^{\text{bg}})$, a subject latent $\bm{z}^{\text{subj}}=\text{VAE}(\bm{x}^{\text{subj}})$, a inpainted latent $\bm{z}^{\text{init}}=\text{VAE}\left(\text{Inpainting}\left[\bm{x}^{\text{bg}},\bm{M}^{\text{user}},\text{VLM}(\bm{x}^{\text{subj}})\right]\right)$, a user mask $\bm{M}^{\text{user}}$.
 2:Output: The composition latent $\bm{z}_{0}$}.
@@
 3.4 Adaptive Background Blending
-Previous methods typically rely on the user-provided mask $\bm{M}^{\text{user}}$ to preserve the background during each denoising step, blending as $\bm{z}_{t}=\bm{M}^{\text{user}}\odot\bm{z}_{t}+(1-\bm{M}^{\text{user}})\odot\bm{z}^{\text{bg}}_{t}$, but this often introduces visible seams along mask boundaries (see the first row of Fig.
+Previous methods typically rely on the user-provided mask $\bm{M}^{\text{user}}$ to preserve the background during each denoising step, blending as $\bm{z}^{\prime}_{t}=\bm{M}^{\text{user}}\odot\bm{z}_{t}+(1-\bm{M}^{\text{user}})\odot\bm{z}^{\text{bg}}_{t}$, but this often introduces visible seams along mask boundaries (see the first row of Fig.
 5).
@@
 Through empirical evaluation, we find that leveraging $\bm{M}^{\text{attn}}$ during the initial denoising steps $(t>\tau)$ sufficiently mitigates visible seams along mask boundaries, ensuring high-fidelity scene coherence.
-Figure 6: Qualitative comparison of our method with multiple baselines in challenging scenarios, drawn from our benchmark dataset. More qualitative comparisons are available in Appendix K.
+Figure 6: Qualitative comparison of our method with multiple baselines in challenging scenarios, drawn from our benchmark dataset. More qualitative comparisons are available in Appendix O.
 4 Experiments
@@
 Metrics.
-Previous methods primarily adapt CLIP-I (Radford et al., 2021) and DINOv2 (Oquab et al., 2023) to assess subject identity consistency.
+Previous methods primarily adapt CLIP-I (Radford et al., 2021) and DINOv2 (Oquab et al., 2024) to assess subject identity consistency.
 However, these features capture high-level semantic information that may not fully align with human perception of finer-grained attributes.
@@
 An analysis of identity consistency metrics is provided in Appendix I.
-To evaluate overall image quality, we use ImageReward (IR) (Xu et al., 2023) and VisionReward (VR) (Xu et al., 2024), which are fine-grained, multi-dimensional reward models that more accurately reflect human preferences.
-For background consistency, LPIPS (Zhang et al., 2018) and SSIM (Wang et al., 2004) are employed.
+For overall image quality, we use ImageReward (IR) (Xu et al., 2023) and VisionReward (VR) (Xu et al., 2024), fine-grained reward models that more accurately reflect human preferences.
+To more comprehensively evaluate composition quality, we further include three UnifiedReward variants (Wang et al., 2025c; d) and HPSv3 (Ma et al., 2025b) in Appendix L.
+Background consistency is measured using LPIPS (Zhang et al., 2018) and SSIM (Wang et al., 2004).
 Implementation Details.
 In our experiment, we used FLUX.1-dev, a 12B-parameter flow matching model, as the base model, combined with InstantCharacter (Tao et al., 2025) as the adapter (Additional results on SDXL, SD3.5, and PixArt are presented in Appendix E).
-Our approach also supports per-concept LoRA (Hu et al., 2021a), which requires test-time tuning (Ruiz et al., 2023b) but delivers superior identity consistency compared to an open-domain adapter, making it ideal for scenarios demanding precise identity preservation.
-The denoising process consists of 20 steps, with the inpainted image perturbed to timestep 15 and denoising initiated from that point.
+Our approach also supports per-concept LoRA (Hu et al., ), which requires test-time tuning (Ruiz et al., 2023) but delivers superior identity consistency compared to an open-domain adapter, making it ideal for scenarios demanding precise identity preservation.
+The denoising schedule consists of 20 steps, with the inpainted image perturbed to timestep 15 and denoising initiated from that point.
+We use Flux.1 Fill (Black Forest Labs, 2024b) as the inpainting model and BLIP-3 (Xue et al., 2024) as the VLM.
 Additional details and hyperparameters are provided in Appendix J.
@@
 Metrics shown in pink are those specifically trained to better align with human preferences.
-Abbreviations: IRF: Instance Retrieval Features; IR = ImageReward; VR = VisionReward.
-Bench Method
-Training
--Free
-Base
-Model
-External
-Model
-Subject Identity Consistency Background Image Quality
-CLIP-I $\uparrow$ DINOv2 $\uparrow$ IRF $\uparrow$ DreamSim $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ IR $\uparrow$ VR $\uparrow$
-Dream-
-Edit-
-Bench
-MADD (He et al., 2024) SD DINO 0.7118 0.6279 0.4333 0.5810 0.0604 0.8182 -0.2545 2.7011
-ObjectStitch (Song et al., 2023) SD VIT 0.7567 0.6930 0.5525 0.5093 0.0190 0.8316 0.0791 3.2416
-DreamCom (Lu et al., 2023c) SD LoRA 0.7414 0.6749 0.5597 0.5626 0.0200 0.8283 0.1873 3.5053
-AnyDoor (Chen et al., 2024c) SD DINO 0.8183 0.7283 0.7714 0.3764 0.0251 0.8894 0.4511 3.3946
-UniCombine (Wang et al., 2025) FLUX LoRA 0.8058 0.7332 0.7579 0.3984 0.0050 0.9397 0.4565 3.6108
-PBE (Yang et al., 2023) SD - 0.7742 0.7040 0.5845 0.4985 0.0197 0.8287 0.2083 3.3482
-TIGIC (Li et al., 2024b) SD - 0.7226 0.6718 0.4711 0.6108 0.0584 0.8153 -0.1332 2.9873
-TALE (Pham et al., 2024) SD - 0.7329 0.6604 0.5007 0.6176 0.0392 0.8251 -0.1502 3.1349
-TF-ICON (Lu et al., 2023d) SD - 0.7479 0.6865 0.5179 0.5441 0.0582 0.8111 0.0816 3.2823
-DreamEdit (Li et al., 2023b) SD LoRA, VIT 0.7703 0.7151 0.6147 0.5047 0.0140 0.9775 0.1744 3.1775
-EEdit (Yan et al., 2025) FLUX - 0.6998 0.6590 0.4438 0.6160 0.0039 0.9475 0.0216 3.3606
-Ours-Adapter FLUX Adapter 0.8086 0.7415 0.7702 0.3730 0.0236 0.8959 0.5709 3.6234
-Ours-LoRA FLUX LoRA 0.8125 0.7452 0.7900 0.3577 0.0271 0.8847 0.5906 3.6161
-Complex-
-Compo
-MADD (He et al., 2024) SD DINO 0.6780 0.5993 0.3638 0.5979 0.0781 0.5658 -0.0088 2.6582
-ObjectStitch (Song et al., 2023) SD VIT 0.7608 0.7077 0.5513 0.4717 0.0388 0.6357 0.2482 3.4411
-DreamCom (Lu et al., 2023c) SD LoRA 0.648 0.5692 0.2788 0.8192 0.0389 0.6342 -0.0778 3.4409
-AnyDoor (Chen et al., 2024c) SD DINO 0.7982 0.7052 0.7319 0.4493 0.0299 0.7262 0.3804 3.3787
-UniCombine (Wang et al., 2025) FLUX LoRA 0.7361 0.6552 0.5380 0.5682 0.0237 0.7077 0.2470 3.5454
-PBE (Yang et al., 2023) SD - 0.7537 0.6802 0.5189 0.5187 0.0397 0.6321 0.2139 3.4310
-TIGIC (Li et al., 2024b) SD - 0.6913 0.6329 0.3848 0.6549 0.0929 0.6228 -0.131 2.8898
-TALE (Pham et al., 2024) SD - 0.6816 0.6151 0.3799 0.6773 0.059 0.6334 0.0783 3.4498
-TF-ICON (Lu et al., 2023d) SD - 0.6987 0.6435 0.4167 0.6030 0.0815 0.6216 0.1798 3.4323
-DreamEdit (Li et al., 2023b) SD LoRA, VIT 0.7314 0.6722 0.5069 0.5670 0.0468 0.7201 0.1212 3.2531
-EEdit (Yan et al., 2025) FLUX - 0.6713 0.6153 0.3797 0.6821 0.0226 0.7107 0.1433 3.5009
-Ours-Adapter FLUX Adapter 0.7721 0.7107 0.6764 0.4294 0.0404 0.7789 0.4090 3.6020
-Ours-LoRA FLUX LoRA 0.7999 0.7384 0.7659 0.3542 0.0430 0.7634 0.4246 3.5951
+Abbreviations: IRF= Instance Retrieval Features; IR = ImageReward; VR = VisionReward; URE = UnifiedReward-Edit-qwen3vl-8b.
+Bench Method Training -Free Base Model External Model Subject Identity Consistency Background Image Quality
+CLIP-I $\uparrow$ DINOv2 $\uparrow$ IRF $\uparrow$ DreamSim $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ IR $\uparrow$ VR $\uparrow$ HPS $\uparrow$ URE $\uparrow$
+Dream- Edit- Bench (220) Flux.1 Fill (Black Forest Labs, 2024b) FLUX - 0.7328 0.6745 0.5754 0.5233 0.0166 0.9076 0.5577 3.5997 8.6432 21.5812
+MADD (He et al., 2024) SD DINO 0.7118 0.6279 0.4333 0.5810 0.0604 0.8182 -0.2545 2.7011 1.2443 13.8148
+ObjectStitch (Song et al., 2023) SD VIT 0.7567 0.6930 0.5525 0.5093 0.0190 0.8316 0.0791 3.2416 7.4529 19.1886
+DreamCom (Lu et al., 2023c) SD LoRA 0.7414 0.6749 0.5597 0.5626 0.0200 0.8283 0.1873 3.5053 5.9324 19.9296
+AnyDoor (Chen et al., 2024c) SD DINO 0.8183 0.7283 0.7714 0.3764 0.0251 0.8894 0.4511 3.3946 8.4867 19.0989
+UniCombine (Wang et al., 2025a) FLUX LoRA 0.8058 0.7332 0.7579 0.3984 0.0050 0.9397 0.4565 3.6108 8.8415 21.7080
+PBE (Yang et al., 2023) SD - 0.7742 0.7040 0.5845 0.4985 0.0197 0.8287 0.2083 3.3482 8.3789 20.2137
+TIGIC (Li et al., 2024b) SD - 0.7226 0.6718 0.4711 0.6108 0.0584 0.8153 -0.1332 2.9873 5.2676 17.1000
+TALE (Pham et al., 2024) SD - 0.7329 0.6604 0.5007 0.6176 0.0392 0.8251 -0.1502 3.1349 6.3773 18.0784
+TF-ICON (Lu et al., 2023d) SD - 0.7479 0.6865 0.5179 0.5441 0.0582 0.8111 0.0816 3.2823 7.2643 18.2716
+DreamEdit (Li et al., 2023b) SD LoRA, VIT 0.7703 0.7151 0.6147 0.5047 0.0140 0.9775 0.1744 3.1775 6.0250 15.7636
+EEdit (Yan et al., 2025) FLUX - 0.6998 0.6590 0.4438 0.6160 0.0039 0.9475 0.0216 3.3606 6.6689 19.5603
+Ours-Adapter FLUX Adapter 0.8086 0.7415 0.7702 0.3730 0.0236 0.8959 0.5709 3.6234 8.8861 22.0182
+Ours-LoRA FLUX LoRA 0.8125 0.7452 0.7900 0.3577 0.0271 0.8847 0.5906 3.6161 8.8688 21.9421
+Complex- Compo (300) Flux.1 Fill (Black Forest Labs, 2024b) FLUX - 0.7108 0.6475 0.5466 0.6018 0.0232 0.7442 0.4088 3.5737 8.7376 19.7712
+MADD (He et al., 2024) SD DINO 0.6780 0.5993 0.3638 0.5979 0.0781 0.5658 -0.0088 2.6582 5.9673 13.0567
+ObjectStitch (Song et al., 2023) SD VIT 0.7608 0.7077 0.5513 0.4717 0.0388 0.6357 0.2482 3.4411 8.8389 18.8283
+DreamCom (Lu et al., 2023c) SD LoRA 0.648 0.5692 0.2788 0.8192 0.0389 0.6342 -0.0778 3.4409 7.9884 18.6143
+AnyDoor (Chen et al., 2024c) SD DINO 0.7982 0.7052 0.7319 0.4493 0.0299 0.7262 0.3804 3.3787 8.9760 18.3550
+UniCombine (Wang et al., 2025a) FLUX LoRA 0.7361 0.6552 0.5380 0.5682 0.0237 0.7077 0.2470 3.5454 8.8999 19.8529
+PBE (Yang et al., 2023) SD - 0.7537 0.6802 0.5189 0.5187 0.0397 0.6321 0.2139 3.4310 8.5923 18.9507
+TIGIC (Li et al., 2024b) SD - 0.6913 0.6329 0.3848 0.6549 0.0929 0.6228 -0.131 2.8898 7.6630 16.4301
+TALE (Pham et al., 2024) SD - 0.6816 0.6151 0.3799 0.6773 0.059 0.6334 0.0783 3.4498 8.7351 18.7567
+TF-ICON (Lu et al., 2023d) SD - 0.6987 0.6435 0.4167 0.6030 0.0815 0.6216 0.1798 3.4323 9.3258 18.2366
+DreamEdit (Li et al., 2023b) SD LoRA, VIT 0.7314 0.6722 0.5069 0.5670 0.0468 0.7201 0.1212 3.2531 8.0434 15.8934
+EEdit (Yan et al., 2025) FLUX - 0.6713 0.6153 0.3797 0.6821 0.0226 0.7107 0.1433 3.5009 8.7835 19.7348
+Ours-Adapter FLUX Adapter 0.7721 0.7107 0.6764 0.4294 0.0404 0.7789 0.4090 3.6020 9.6485 20.7349
+Ours-LoRA FLUX LoRA 0.7999 0.7384 0.7659 0.3542 0.0430 0.7634 0.4246 3.5951 9.8418 21.0326
 4.2 Experimental Results
-We compare our method with two categories of baselines: (1) Training-based methods (6 in total): UniCombine (Wang et al., 2025), AnyDoor (Chen et al., 2024c), Paint by Example (PBE) (Yang et al., 2023), ObjectStitch (Song et al., 2023), MADD (He et al., 2024), and DreamCom (Lu et al., 2023c); (2) Training-free methods (5 in total): EEdit (Yan et al., 2025), TIGIC (Li et al., 2024b), DreamEdit (Li et al., 2023b), TF-ICON (Lu et al., 2023d), and TALE (Pham et al., 2024).
+We compare our method with two categories of baselines: (1) Training-based methods (6 in total): UniCombine (Wang et al., 2025a), AnyDoor (Chen et al., 2024c), Paint by Example (PBE) (Yang et al., 2023), ObjectStitch (Song et al., 2023), MADD (He et al., 2024), and DreamCom (Lu et al., 2023c); (2) Training-free methods (5 in total): EEdit (Yan et al., 2025), TIGIC (Li et al., 2024b), DreamEdit (Li et al., 2023b), TF-ICON (Lu et al., 2023d), and TALE (Pham et al., 2024).
 As shown in Tab.
@@
 4.3 Ablation Study
-Figure 7: Qualitative ablation study comparing different variants of our framework.
 Table 2: Ablation study examining the impact of key components on DreamEditBench.
-Method
-MSA
-DSG
-ABB
-Subject Identity Consistency Background Image Quality
+Method MSA DSG ABB Subject Identity Consistency Background Image Quality
 CLIP-I DINOv2 IRF DreamSim LPIPS SSIM IR VR
@@
 5, 7), it is less well captured by quantitative metrics, since LPIPS and SSIM primarily assess structural similarity rather than perceptual smoothness.
+Figure 7: Qualitative ablation study comparing different variants of our framework.
 5 Conclusion
@@
 While current customization adapters already perform well, the potential of our method will continue to improve as advancements are made in the field of open-domain customization adapters.
+Ethics Statement
+Our framework provides an accessible way for people without professional artistic skills to create image compositions.
+While this technology offers significant benefits, it also carries the risk of misuse for malicious purposes, such as harassment or spreading misinformation.
+Additionally, our framework relies on pretrained large-scale T2I models, which may inadvertently introduce social and cultural biases.
+Therefore, using these models raises ethical concerns and requires careful consideration.
+We therefore urge users to exercise caution and use this tool responsibly for appropriate purposes only.
+Reproducibility Statement
+We have taken several measures to ensure the reproducibility of our results.
+Algorithm 1 presents the pseudocode of our method.
+Details of the implementation for the main experiments are provided in Sec.
+4.1, while the hyperparameter configurations are listed in Appendix J.
+The source code and the ComplexCompo dataset will be released publicly.
+This work adheres to the reproducibility requirements set by ICLR.
+Acknowledgments
+This research is supported by the National Research Foundation, Singapore and Infocomm Media Development Authority under its Trust Tech Funding Initiative.
+Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore and Infocomm Media Development Authority.
 Appendix
@@
 Training-based methods fine-tune foundational models using datasets tailored for image composition.
-Early methods like Paint by Example (Yang et al., 2022) and ObjectStitch (Song et al., 2022) use CLIP to encode subject features, ensuring high semantic similarity between inserted objects and reference images.
+Early methods like Paint by Example (Yang et al., 2023) and ObjectStitch (Song et al., 2023) use CLIP to encode subject features, ensuring high semantic similarity between inserted objects and reference images.
 These approaches use image augmentation to create training datasets, enabling effective training.
@@
 DreamCom (Lu et al., 2023c) and MureObjectStitch (Chen et al., 2024a) fine-tune models with small sets of reference images to preserve subject identity.
-AnyDoor (Chen et al., 2024c), IMPRINT (Song et al., 2024b), and E-MD3C (Pham et al., 2025) leverage DINOv2 to enhance identity fidelity and control over shape and pose, drawing supervision from video data.
+AnyDoor (Chen et al., 2024c), IMPRINT (Song et al., 2024), and E-MD3C (Pham et al., 2025) leverage DINOv2 to enhance identity fidelity and control over shape and pose, drawing supervision from video data.
 MimicBrush (Chen et al., 2024b) similarly uses video-derived supervision for imitative editing.
-In contrast, MADD (He et al., 2024), ObjectMate (Winter et al., 2024), and OmniPaint (Yu et al., 2025b) employ image inpainting models to generate higher-quality training datasets compared to those based on image or video augmentation.
+In contrast, MADD (He et al., 2024), ObjectMate (Winter et al., 2025), and OmniPaint (Yu et al., 2025c) employ image inpainting models to generate higher-quality training datasets compared to those based on image or video augmentation.
 Multitwine (Tarrés et al., 2025) enables the integration of multiple objects, capturing interactions from simple positional relationships to complex actions requiring reposing.
 DreamFuse (Huang et al., 2025) uses a Positional Affine mechanism to embed foreground size and position into the background, fostering effective foreground-background interaction through shared attention.
-Insert Anything (Song et al., 2025) and UniCombine (Wang et al., 2025) introduces a FLUX-based, multi-conditional generative framework that handles diverse condition combinations.
+Insert Anything (Song et al., 2025a) and UniCombine (Wang et al., 2025a) introduces a FLUX-based, multi-conditional generative framework that handles diverse condition combinations.
 However, these methods often bias the generative priors of base models toward curated datasets, resulting in unnatural compositions, such as implausible object-environment interactions (e.g., missing or unrealistic shadows and reflections).
@@
 TIGIC (Li et al., 2024b) focuses on preserving non-target areas during composition.
-Thinking Outside the BBox (Tarrés et al., 2024) enables unconstrained image compositing, unbound by input masks.
+Thinking Outside the BBox (Canet Tarrés et al., 2024) enables unconstrained image compositing, unbound by input masks.
 FreeCompose (Chen et al., 2024e) employs a pipeline of object removal, image harmonization, and semantic composition.
-DreamEdit (Li et al., 2023b), UniCanvas (Jin et al., 2025), and Magic Insert (Ruiz et al., 2024) use test-time tuning to fine-tune models during inference.
-Add-it (Tewel et al., 2024) enables text-guided object insertion on FLUX, where users describe objects via text prompts instead of reference images.
+DreamEdit (Li et al., 2023b), UniCanvas (Jin et al., 2025), and Magic Insert (Ruiz et al., 2025) use test-time tuning to fine-tune models during inference.
+Add-it (Tewel et al., ) enables text-guided object insertion on FLUX, where users describe objects via text prompts instead of reference images.
 EEdit (Yan et al., 2025) recently improves TF-ICON on FLUX, introducing step-skipping to reduce time costs and spatial locality caching to minimize redundancy.
@@
 A.2 General Image Editing
-Recent and significant advancements in text-to-image generative models have enhanced numerous applications (Zhou et al., 2025c; a; b; 2023; 2024a; 2024b; 2024c; Zhao et al., 2024a; d; c; 2025b; 2025a; b; Xie et al., 2024; Wang et al., 2023b; Wang & Chen, 2025a; b; Lu et al., 2024b; Li et al., 2025; Gao et al., 2024; 2025; Zhu et al., 2024; Yu et al., 2025a).
+Recent and significant advancements in text-to-image generative models have enhanced numerous applications (Zhou et al., 2023; 2024a; 2024b; 2024c; 2025a; 2025b; Wang et al., 2025b; Wang & Chen, 2025a; b; Zhao et al., 2024a; 2025b; 2025a; b; 2026; Li et al., 2025; Gao et al., 2024; 2025; Zhu et al., 2024; Lu et al., 2024b; Ren et al., 2025; Yu et al., 2025b).
 Instruction-based image editing has evolved rapidly.
-Early systems relied on modular, two-stage pipelines: a multimodal language model first produced textual prompts, spatial guidance, or synthetic instruction–image pairs, and a separate diffusion model then executed the edit—as in InstructEdit (Wang et al., 2023a), InstructPix2Pix (Brooks et al., 2022), MagicBrush (Zhang et al., 2023b), and BrushEdit (Li et al., 2024c).
+Early systems relied on modular, two-stage pipelines: a multimodal language model first produced textual prompts, spatial guidance, or synthetic instruction–image pairs, and a separate diffusion model then executed the edit—as in InstructEdit (Wang et al., 2023), InstructPix2Pix (Brooks et al., 2023), MagicBrush (Zhang et al., 2023b), and BrushEdit (Li et al., 2024c).
 Recent work has shifted toward tightly integrated, instruction-centric architectures.
-Models such as SmartEdit (Huang et al., 2024), X2I (Ma et al., 2025), RPG (Yang et al., 2024), AnyEdit (Yu et al., 2024), and UltraEdit (Zhao et al., 2024e) embed routing, task-aware objectives, and fine-grained controls directly into the network, yielding higher fidelity and more precise manipulation.
-Unified generation-and-editing frameworks (e.g., OmniGen (Xiao et al., 2024), ACE (Han et al., 2024), ACE++ (Mao et al., 2025), Lumina-OmniLV (Pu et al., 2025), Qwen2VL-Flux (Lu, 2024), DreamEngine (Chen et al., 2025), MetaQueries (Pan et al., 2025), Hidream-E1 (HiDream-ai, 2025)) treat editing as one capability of an end-to-end vision-language model, often fusing language embeddings with diffusion latents to provide context-aware, pixel-level control.
+Models such as SmartEdit (Huang et al., 2024), X2I (Ma et al., 2025a), RPG (Yang et al., 2024), AnyEdit (Yu et al., 2025a), and UltraEdit (Zhao et al., 2024c) embed routing, task-aware objectives, and fine-grained controls directly into the network, yielding higher fidelity and more precise manipulation.
+Unified generation-and-editing frameworks (e.g., OmniGen (Xiao et al., 2025), ACE (Han et al., ), ACE++ (Mao et al., 2025), Lumina-OmniLV (Pu et al., 2025), Qwen2VL-Flux (Lu, 2024), DreamEngine (Chen et al., 2025), MetaQueries (Pan et al., 2025), Hidream-E1 (HiDream-ai, 2025)) treat editing as one capability of an end-to-end vision-language model, often fusing language embeddings with diffusion latents to provide context-aware, pixel-level control.
 Efficiency has advanced in parallel: ICEdit (Zhang et al., 2025) couples LoRA with mixture-of-experts tuning and optimized noise initialization, while SuperEdit (Ming et al., 2025) relies on higher-quality data and contrastive supervision to sustain performance at lower cost.
@@
 Our framework accommodates both categories, so we provide two corresponding variants in the main paper.
-Test-time fine-tuning methods (Gal et al., 2022a; Ruiz et al., 2023a) adapt a pre-trained T2I model to a small set of reference images (typically 3 to 5 images).
+Test-time fine-tuning methods (Gal et al., 2022; Ruiz et al., 2023) adapt a pre-trained T2I model to a small set of reference images (typically 3 to 5 images).
 Although this step adds computational cost and latency, it offers the greatest flexibility for diverse customization requirements.
 Such methods are commonly grouped into three subclasses: data regularization, weight regularization, and loss regularization.
-In the data-regularization family, DreamBooth (Ruiz et al., 2023a) limits overfitting by generating superclass images with the base T2I model and training on both reference and regularization images; Custom Diffusion (Kumari et al., 2023) improves regularization quality by retrieving real images; and Specialist Diffusion (Lu et al., 2023a) applies extensive data augmentation.
-Weight-regularization approaches (Gal et al., 2022b; Hu et al., 2021b; Han et al., 2023; Qiu et al., 2023) confine updates to carefully chosen parameters, such as the subject-specific text embeddings or the singular values of weight matrices.
-Loss-regularization approaches, including Specialist Diffusion (Lu et al., 2023a), MagiCapture (Hyung et al., 2023), and FaceChain-SuDe (Qiao et al., 2024), introduce objective terms that respectively maximize CLIP-space similarity between generated and reference images, disentangle identity and style via masked facial reconstruction, or encourage correct superclass classification.
+In the data-regularization family, DreamBooth (Ruiz et al., 2023) limits overfitting by generating superclass images with the base T2I model and training on both reference and regularization images; Custom Diffusion (Kumari et al., 2023) improves regularization quality by retrieving real images; and Specialist Diffusion (Lu et al., 2023a) applies extensive data augmentation.
+Weight-regularization approaches (Gal et al., 2022; Hu et al., ; Han et al., 2023; Qiu et al., 2023) confine updates to carefully chosen parameters, such as the subject-specific text embeddings or the singular values of weight matrices.
+Loss-regularization approaches, including Specialist Diffusion (Lu et al., 2023a), MagiCapture (Hyung et al., 2024), and FaceChain-SuDe (Qiao et al., 2024), introduce objective terms that respectively maximize CLIP-space similarity between generated and reference images, disentangle identity and style via masked facial reconstruction, or encourage correct superclass classification.
 Zero-shot image customization methods avoid subject-specific fine-tuning at inference time but rely on extensive pre-training.
 For general subject customization, InstantBooth (Shi et al., 2024a) adds a visual encoder that captures coarse-to-fine image features from the references; BLIP-Diffusion (Li et al., 2024a) fine-tunes BLIP-2 (Li et al., 2023a) to extract multimodal subject representations; ELITE (Wei et al., 2023) maps reference images into hierarchical textual tokens through global and local networks; and Song et al.
-(Song et al., 2024a) enhance textual control by removing the projection of visual embeddings onto textual embeddings.
+(Song et al., 2025b) enhance textual control by removing the projection of visual embeddings onto textual embeddings.
 For facial customization, InstantID (Wang et al., 2024b) isolates facial regions from reference images to extract appearance and structural cues.
-For style customization, InstantStyle (Wang et al., 2024a) identifies style-controlling layers and injects IP-Adapter features (Ye et al., 2023a) into those layers to achieve style transfer.
-InstantCharacter (Tao et al., 2025), IP-Adapter (Ye et al., 2023b), and PuLID (Guo et al., 2024) have each released versions compatible with the FLUX model.
+For style customization, InstantStyle (Wang et al., 2024a) identifies style-controlling layers and injects IP-Adapter features (Ye et al., 2023) into those layers to achieve style transfer.
+InstantCharacter (Tao et al., 2025), IP-Adapter (Ye et al., 2023), and PuLID (Guo et al., 2024) have each released versions compatible with the FLUX model.
 Appendix B Visualizing the Impact of Adaptive Background Blending
@@
 A similar argument applies to the value matrix $\bm{V}$, as the output of self-attention, $(\bm{QK}^{\mathsf{T}})\bm{V}$, involves $\bm{V}$ in a post-multiplication step, and convolution on $\bm{V}$ does not commute with the attention weight computation in the same manner.
+C.3 Implementation Details of Gaussian Blurring
+For the 2D Gaussian filtering step, we adopt the implementation provided by the kornia library.
+Following standard engineering practice, we set the kernel radius to $r=3\sigma$, since three standard deviations capture approximately $99.7\%$ of the Gaussian mass.
+Consequently, the kernel size is chosen as the nearest odd integer to $2r=6\sigma$.
+In all of our experiments we use $\sigma=10$ (see Tab.
+5).
+The procedure is applied to query embeddings by reshaping them into spatial feature maps, performing Gaussian smoothing, and then mapping them back into the sequence domain prior to attention computation.
+The full workflow is given below:
+Algorithm 2 Implementation Details of 2D Gaussian Blurring
+1:Input: A query matrix $\mathbf{Q}\in\mathbb{R}^{B\times L\times D}$, key $\mathbf{K}$, value $\mathbf{V}$, spatial dimensions $(H,W)$, Gaussian standard deviation $\sigma$.
+2:Output: Attention output $\mathbf{O}$.
+3:// Reshape query into spatial tensor
+[…]
+[…]
+6:// Construct Gaussian kernel size
+[…]
+[…]
+[…]
+[…]
+11:// Apply 2D Gaussian smoothing
+[…]
+13:// Reshape smoothed queries back to sequence form
+[…]
+[…]
+16:// Compute attention
+[…]
+[…]
+19:return $\mathbf{O}$
 Appendix D Evaluating Cross-Attention Map Accuracy via IoU
@@
 In the main text, we showcase the effectiveness of our approach on Flux, a leading open-source model.
-To establish its generalizability, we conduct experiments on SDXL (Podell et al., 2024), SD3.5 (Esser et al., 2024), and PixArt (Chen et al., 2023) across DreamEditBench and ComplexCompo.
+To establish its generalizability, we conduct experiments on SDXL (Podell et al., 2024), SD3.5 (Esser et al., 2024), and PixArt (Chen et al., ) across DreamEditBench and ComplexCompo.
 The results are presented in Tab.
@@
 Abbreviations: IRF: Instance Retrieval Features; IR = ImageReward; VR = VisionReward.
-Bench Method
-Base
-Model
-Subject Identity Consistency Background Image Quality
+Bench Method Base Model Subject Identity Consistency Background Image Quality
 CLIP-I $\uparrow$ DINOv2 $\uparrow$ IRF $\uparrow$ DreamSim $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ IR $\uparrow$ VR $\uparrow$
-DreamEdit-
-Bench
+DreamEdit- Bench (220) Flux.1 Fill FLUX 0.7328 0.6745 0.5754 0.5233 0.0166 0.9076 0.5577 3.5997
 Ours-Adapter SDXL 0.7944 0.7334 0.7659 0.3761 0.0238 0.8922 0.5621 3.6158
@@
 Ours-LoRA FLUX 0.8125 0.7452 0.7900 0.3577 0.0271 0.8847 0.5906 3.6161
-Complex-
-Compo
+Complex- Compo (300) Flux.1 Fill FLUX 0.7108 0.6475 0.5466 0.6018 0.0232 0.7442 0.4088 3.5737
 Ours-Adapter SDXL 0.7657 0.7084 0.6862 0.4457 0.0457 0.7612 0.3894 3.5987
@@
 Our benchmark consists of 300 triplets, each comprising a subject image, a background image, and a bounding box.
-The subject images are identical to those used in DreamEditBench (Li et al., 2023b; Ruiz et al., 2023a), while the background images are sampled from the OpenImages dataset (Kuznetsova et al., 2020).
+The subject images are identical to those used in DreamEditBench (Li et al., 2023b; Ruiz et al., 2023), while the background images are sampled from the OpenImages dataset (Kuznetsova et al., 2020).
 These backgrounds exhibit a variety of aspect ratios and resolutions, including landscape and portrait formats, such as $768\times 1088$, $768\times 1072$, $768\times 1024$, $768\times 1152$, $024\times 768, 11$$152\times 768$, $1200\times 768$, $848\times 768$, and $1360\times 768$.
@@
 Appendix I Subject Identity Metrics Analysis
-In our experiments we found that widely-used subject-identity metrics such as CLIP-I (Radford et al., 2021) and DINOv2 (Oquab et al., 2023) correlate poorly with human preferences.
+In our experiments we found that widely-used subject-identity metrics such as CLIP-I (Radford et al., 2021) and DINOv2 (Oquab et al., 2024) correlate poorly with human preferences.
 Because they focus almost exclusively on semantic similarity, they ignore appearance changes introduced by lighting, shadows, reflections, and surrounding context.
 Fig.
-12(b) presents several image pairs produced by AnyDoor (left) and by our method (right); the corresponding CLIP-I $(\uparrow)$ (Radford et al., 2021), DINOv2 $(\uparrow)$ (Oquab et al., 2023), IRF $(\uparrow)$ (Shao & Cui, 2022), and DreamSim $(\downarrow)$ (Fu et al., 2023) scores are shown beneath each image, with the better score highlighted in red.
+12(b) presents several image pairs produced by AnyDoor (left) and by our method (right); the corresponding CLIP-I $(\uparrow)$ (Radford et al., 2021), DINOv2 $(\uparrow)$ (Oquab et al., 2024), IRF $(\uparrow)$ (Shao & Cui, 2022), and DreamSim $(\downarrow)$ (Fu et al., 2023) scores are shown beneath each image, with the better score highlighted in red.
 Although the AnyDoor results are visibly less realistic and less consistent, they nevertheless receive higher CLIP-I and DINOv2 scores, and in most cases higher IRF scores, demonstrating that these measures do not faithfully capture compositional quality.
@@
 5.
-Each baseline is implemented according to the configuration settings recommended in its original publication.
-The repositories utilized for each baseline are listed below:
+Under Denoising Setup, “Total steps” refers to the full diffusion/noising schedule, which specifies a sequence of 20 values of $\sigma_{t}$ across timesteps $t$.
+However, our method does not begin denoising at the first timestep.
+As shown in Algorithm 1, it starts at $t_{1}=14$ (Ours-Adapter), resulting in 15 denoising steps in total.
+Under MSA loss, “Step Range” indicates the subset of denoising steps to which MSA optimization is applied.
+For the adapter setting, the MSA loss is applied only to the first three denoising steps (from $t=14$ to $t=12$).
+In addition, since a LoRA is trained for a specific subject, it provides a more tailored prior than the generic adapter, allowing it to converge in fewer steps and with lower overall compute.
+Each baseline is implemented according to the configuration settings recommended in its original publication. The repositories utilized for each baseline are listed below:
 Table 5: Hyperparameters of Our Frameworks. Bin Thresh = Binarization Threshold; #iter = Number of Iterations.
 Variant Denoising Setup Manifold-Steered Anchor Loss Degradation-Suppression Guidance Adaptive Background Blending
-Total steps
-$(T-1)\to 0$
-Start
-step $t_{1}$
-Step Range
-$t_{1}\to\tau$
-Learning Rate
-$\alpha$
-#iters
-$k$
-Step Range
-$t_{1}\to 0$
-Scale
-$\eta$
-Blur
-$\sigma$
-Step Range
-$(t_{1}-1)\to 1$
-Bin
-Thresh
-Dilation
-Kernel Size
+Total steps $(T-1)\to 0$ Start step $t_{1}$ Step Range $t_{1}\to\tau$ Learning Rate $\alpha$ #iters $k$ Step Range $t_{1}\to 0$ Scale $\eta$ Blur $\sigma$ Step Range $(t_{1}-1)\to 1$ Bin Thresh Dilation Kernel Size
 Ours-Adapter $19\to 0$ 14 $14\to 12$ 500, 750, 1000 10 $14\to 0$ 0.5 10 $13\to 1$ 0.2 3
 Ours-LoRA $19\to 0$ 13 $13\to 12$ 50, 300 2 $13\to 0$ 0.7 10 $12\to 1$ 0.4 3
-Appendix K Additional Qualitative Results
+Appendix K Discussion on Inversion vs. One-Step Forward Diffusion
+We provide an expanded discussion of our design choice between inversion and one-step forward diffusion to better clarify the motivation behind our approach.
+In training-free image editing, both inversion and one-step forward diffusion are commonly used to obtain a noisy latent that serves as the starting point for subsequent optimization or denoising.
+In our framework, we intentionally adopt one-step forward diffusion as a practical substitute for inversion.
+Our method does not depend on the initial latent to preserve object identity.
+Instead, the MSA loss extracts object-specific information from the adapter/LoRA and injects it into the latent during the editing process.
+This design choice is motivated by a practical observation: many modern models are distilled for speed, making accurate inversion difficult to achieve in practice.
+When inversion cannot reliably encode object identity, its benefit becomes limited.
+In such cases, a noisy latent that still retains enough background structure is sufficient as a starting point.
+For this reason, we adopt one-step forward diffusion as a pragmatic replacement for inversion, rather than a theoretically equivalent alternative.
+It offers two advantages: (i) it avoids the accuracy limitations of inversion on distilled models, making it more broadly applicable, and (ii) it is faster than performing an inversion.
+EEdit also proposes an elegant strategy, termed inversion skipping, to accelerate the initialization process.
+In EEdit, the initial latent is obtained via an inversion procedure involving model predictions, while inversion skipping significantly reduces the number of required steps.
+In contrast, our one-step forward diffusion does not perform inversion at all.
+The initial latent is produced by directly sampling noise and adding it to the clean latent at a chosen timestep, without any model prediction.
+This makes our initialization computationally lighter.
+Because FLUX-Dev is a CFG-distilled model, its inversions are relatively imprecise, which we believe contributes to the weaker subject-identity preservation observed in EEdit.
+Our method is therefore designed to avoid reliance on accurate inversion in such distilled settings.
+Appendix L Additional Image Quality Evaluation Using UnifiedReward and HPSv3
+To provide a more comprehensive assessment of composition quality, we further evaluate the methods using three variants of UnifiedReward (Wang et al., 2025c; d) (UnifiedReward-2.0-qwen3vl-8b, UnifiedReward-Edit-qwen3vl-8b, and UnifiedReward-Think-qwen-7b), as well as HPSv3 (Ma et al., 2025b).
+The results are summarized in Tab.
+6.
+Table 6: Image Quality Evaluation Results using HPSv3 and UnifiedReward variants.
+Bench Method HPSv3 UnifiedReward-2.0-qwen3vl-8b UnifiedReward-Edit-qwen3vl-8b UnifiedReward-Think-qwen-7b
+Instruction Quality Average Instruction Quality Average Instruction Quality Average
+Success Overedit Natural Artifact Success Overedit Natural Artifact Success Overedit Natural Artifact
+Dream- Edit- Bench (220) MADD (He et al., 2024) 1.2443 14.2273 12.3727 16.6153 18.6807 15.4740 11.8091 17.8727 13.1045 12.4727 13.8148 14.7553 16.2128 19.1103 20.5724 17.6627
+ObjectStitch (Song et al., 2023) 7.4529 19.9954 16.8721 21.0455 21.8864 19.9499 16.2500 21.2000 18.6545 20.6500 19.1886 20.7500 17.7614 21.0563 20.7676 20.0838
+DreamCom (Lu et al., 2023c) 5.9324 18.1818 15.4818 23.5388 24.0502 20.3132 14.0727 22.5409 21.0091 22.0955 19.9296 18.5263 15.4737 21.6641 21.6953 19.3399
+AnyDoor (Chen et al., 2024c) 8.4867 23.2146 17.8128 20.9772 21.9361 20.9852 19.0909 20.6000 17.4182 19.2864 19.0989 21.2083 18.5694 21.9214 21.5929 20.8230
+UniCombine (Wang et al., 2025a) 8.8415 22.4545 17.8000 23.1682 23.9000 21.8307 20.1591 22.7773 21.2500 22.6455 21.7080 23.0595 16.2381 21.9520 20.8400 20.5224
+PBE (Yang et al., 2023) 8.3789 22.4292 17.5205 22.2773 23.2273 21.3636 18.2318 21.8182 19.5955 21.2091 20.2137 24.1481 17.8025 21.5786 21.9500 21.3698
+TIGIC (Li et al., 2024b) 5.2676 17.3136 14.8045 19.0318 20.3636 17.8784 13.7045 19.4636 17.0864 18.1455 17.1000 17.7294 17.0000 21.0897 20.8828 19.1755
+TALE (Pham et al., 2024) 6.3773 19.6027 15.9863 20.4455 21.4773 19.3780 14.8318 21.6182 17.3227 18.5409 18.0784 21.6071 17.6548 21.0667 20.6933 20.2555
+TF-ICON (Lu et al., 2023d) 7.2643 20.4490 16.7538 20.8273 21.7227 19.9382 15.9045 20.1727 17.9909 19.0182 18.2716 20.9870 16.9870 21.2946 20.7984 20.0168
+DreamEdit (Li et al., 2023b) 6.0250 19.9227 16.5409 19.2773 20.5227 19.0659 14.9273 19.1818 13.3909 15.5545 15.7636 20.4714 16.8143 20.3333 20.7153 19.5836
+EEdit (Yan et al., 2025) 6.6689 18.3790 15.3379 22.0636 23.4818 19.8156 14.2091 22.2864 19.9500 21.7955 19.5603 21.1463 18.2561 22.0635 21.2222 20.6720
+Ours-Adapter 8.8861 23.3727 17.0500 23.5727 23.8136 21.9523 21.3364 22.7591 21.3636 22.6136 22.0182 23.2222 18.7407 22.8448 22.0086 21.7041
+Ours-LoRA 8.8688 23.4545 16.8136 23.7727 23.8500 21.9727 21.1273 22.7455 21.3000 22.5955 21.9421 23.7590 18.5904 22.4853 20.6765 21.3778
+Complex- Compo (300) MADD (He et al., 2024) 5.9673 13.8900 11.8167 18.8633 17.1333 15.4258 12.6800 16.7300 12.5167 10.3000 13.0567 17.6538 15.7115 20.2867 19.5533 17.2524
+ObjectStitch (Song et al., 2023) 8.8389 20.7157 13.1773 21.4200 19.4800 18.6983 17.3567 21.8500 17.1333 18.9733 18.8283 19.8304 14.6518 21.3467 19.5400 18.8394
+DreamCom (Lu et al., 2023c) 7.9884 8.2234 9.0756 23.9178 23.4737 16.1726 5.9507 23.7072 22.3618 22.4375 18.6143 17.8810 15.2857 22.5000 20.9737 19.0509
+AnyDoor (Chen et al., 2024c) 8.9760 21.7633 13.1133 20.5567 18.5300 18.4908 18.4967 21.8300 14.9267 18.1667 18.3550 22.6606 17.6422 21.2467 19.5333 19.8876
+UniCombine (Wang et al., 2025a) 8.8999 15.6747 11.2226 23.0878 22.7230 18.1770 13.7399 23.4966 20.7195 21.4554 19.8529 20.3646 15.7500 21.9764 20.2804 19.6449
+PBE (Yang et al., 2023) 8.5923 19.6151 13.1283 21.9243 20.1349 18.7007 16.8947 21.9605 17.3257 19.6217 18.9507 23.5810 14.7048 21.2039 19.6447 19.6170
+TIGIC (Li et al., 2024b) 7.6630 14.6250 11.9899 21.2357 20.0202 16.9677 12.6027 19.5185 16.9192 16.6801 16.4301 16.7168 16.5398 21.5051 20.2862 18.2956
+TALE (Pham et al., 2024) 8.7351 16.9899 11.3826 22.7600 20.7000 17.9581 15.0100 22.3567 19.0333 18.6267 18.7567 17.8317 15.4455 21.9667 19.9767 18.7955
+TF-ICON (Lu et al., 2023d) 9.3258 17.7047 12.6812 21.9463 20.4799 18.2030 15.1477 21.2919 18.1577 18.3490 18.2366 19.9775 16.7072 21.5101 19.7584 19.2380
+DreamEdit (Li et al., 2023b) 8.0434 17.2600 12.0233 18.5669 17.2578 16.2770 14.7300 19.4600 13.7467 15.6367 15.8934 23.3964 15.6396 20.4333 19.5400 18.9805
+EEdit (Yan et al., 2025) 8.7835 14.9500 11.2567 22.8746 21.8152 17.7241 13.4224 23.1485 20.2601 22.1081 19.7348 22.8058 15.6990 22.3498 20.5941 20.2367
+Ours-Adapter 9.6485 22.6162 14.2525 22.5552 21.9064 20.3326 20.4582 22.6421 18.5518 21.2876 20.7349 24.0300 17.4600 22.2074 20.3913 20.9647
+Ours-LoRA 9.8418 22.9532 13.3779 22.9130 21.7525 20.2492 20.8328 22.8261 19.1940 21.2776 21.0326 24.3400 17.8900 22.3838 19.9596 21.1212
+Appendix M Further Analysis on DSG in SD3.5
+We conduct experiments on SD3.5 by replacing our DSG mechanism with standard negative prompting.
+Specifically, in Eqn.
+3, we substitute the negative velocity $\bm{v}^{\text{neg}}_{\bm{\theta}+\Delta\bm{\theta}}$ with the velocity obtained using a variety of commonly used negative prompts.
+The four sets of negative prompts used are:
+  1. 1.
+    distorted, deformed, glitch, artifacts
+  2. 2.
+    undefined shapes, bad anatomy, unnatural pose
+  3. 3.
+    low quality, worst quality, low resolution, blurry, out of focus
+  4. 4.
+    AI artifacts, melted objects, strange textures
+The quantitative results on the DreamEditBench dataset are presented in Tab. 7.
+Table 7: Comparison of negative prompting versus DSG on SD3.5 on the DreamEditBench.
+Method CLIP $\uparrow$ DINO $\uparrow$ IRF $\uparrow$ DreamSim $\downarrow$ LPIPS $\downarrow$ SSIM $\uparrow$ IR $\downarrow$ VR $\uparrow$
+Ours-SD3.5-Adapter (w/ NP 1) 0.7934 0.7311 0.7608 0.3918 0.0331 0.8812 0.5837 3.5714
+Ours-SD3.5-Adapter (w/ NP 2) 0.8029 0.7388 0.7679 0.3781 0.0259 0.8908 0.5726 3.6102
+Ours-SD3.5-Adapter (w/ NP 3) 0.7958 0.7329 0.7623 0.3892 0.0315 0.8839 0.5814 3.5796
+Ours-SD3.5-Adapter (w/ NP 4) 0.7887 0.7281 0.7574 0.3976 0.0368 0.8765 0.5882 3.5548
+Ours-SD3.5-Adapter (w/ DSG) 0.8054 0.7407 0.7699 0.3745 0.0234 0.8937 0.5701 3.6187
+Key Insights
+These results lead to the following key conclusions:
+  1. 1.
+    Performance Sensitivity: The performance of standard negative prompting is sensitive to the specific wording used (e.g., Prompt 2 performs better than Prompt 4).
+This confirms that the effectiveness of heuristic negative prompts relies heavily on manual, often tedious, prompt engineering.
+  2. 2.
+    Adaptive Guidance: Standard negative prompts generate a degradation direction that is decoupled from the specifics of the input image.
+In contrast, DSG constructs an image-specific low-quality direction by blurring the attention component, making it significantly more adaptive, stable, and effective across diverse inputs without requiring any manual prompt tuning.
+DSG consistently outperforms all tested standard negative prompts across all metrics.
+This analysis confirms that DSG provides a superior, more robust, and automated mechanism for degradation suppression compared to traditional negative prompting, even in models like SD3.5 where negative prompts are generally effective.
+Appendix N Runtime Comparison
+The wall-clock runtime at a $512\times 512$ resolution on an H100 GPU is summarized in the table below.
+In our implementation, we applied qint8 quantization to all FLUX-based methods to accelerate inference and reduce GPU memory usage.
+Additionally, Ours-Adapter can also be run on a 24-GB GPU by enabling CPU offloading.
+Table 8: Runtime and memory usage comparison of various image composition methods at a $512\times 512$ resolution.
+Method Training Free Base Model External Model Time (s) Peak Memory (MB)
+MADD (He et al., 2024) SD DINO 45.73 11708
+ObjectStitch (Song et al., 2023) SD VIT 6.63 8268
+DreamCom (Lu et al., 2023c) SD LoRA 9.87 3388
+AnyDoor (Chen et al., 2024c) SD DINO 8.61 18612
+UniCombine (Wang et al., 2025a) FLUX LoRA 11.98 22711
+PBE (Yang et al., 2023) SD - 3.52 10842
+TIGIC (Li et al., 2024b) SD - 10.82 21640
+TALE (Pham et al., 2024) SD - 8.03 23524
+TF-ICON (Lu et al., 2023d) SD - 24.55 20670
+DreamEdit (Li et al., 2023b) SD LoRA, VIT 99.83 19298
+EEdit (Yan et al., 2025) FLUX - 60.31 26546
+Ours-Adapter FLUX Adapter 38.29 32552
+Ours-LoRA FLUX LoRA 18.08 23519
+Appendix O Additional Qualitative Results
 We offer more qualitative assessment results, including visualizations of all baselines, presented in Figs. 13 to 18.
-Appendix L LLM Usage Statement
+Appendix P LLM Usage Statement
 We used large language models for text polishing and grammar correction during manuscript preparation.